### v1

# From Frege to chatGPT: Compositionality in language, cognition, and deep neural networks

Abstract
Compositionality has long been considered a key explanatory property underlying human intelligence: arbitrary concepts can be composed into novel complex combinations, permitting the acquisition of an open ended, potentially infinite expressive capacity from finite learning experiences. Influential arguments have held that neural networks fail to explain this aspect of behavior, leading many to dismiss them as viable models of human cognition. Over the last decade, however, modern deep neural networks (DNNs)—which share the same fundamental design principles as their predecessors—have come to dominate artificial intelligence, exhibiting the most advanced cognitive behaviors ever demonstrated in machines. In particular, large language models (LLMs), DNNs trained to predict the next word on a large corpus of text, have proven capable of sophisticated behaviors such as writing syntactically complex sentences without grammatical errors, producing cogent chains of reasoning, and even writing original computer programs—all behaviors thought to require compositional processing. In this chapter, we survey recent empirical work from machine learning for a broad audience in philosophy, cognitive science, and neuroscience, situating recent breakthroughs in deep learning and LLMs within the broader context of philosophical arguments about compositionality. In particular, our review emphasizes two approaches to endowing neural networks with compositional generalization capabilities: (1) architectural inductive biases, and (2) metalearning, or learning to learn. We also present findings suggesting that LLM pretraining can be understood as a kind of metalearning, and can thereby equip DNNs with compositional generalization abilities in a similar way. We conclude by discussing the implications that these findings may have for the study of compositionality in human cognition and by suggesting avenues for future research.
1 Introduction
Deep neural networks (DNNs) have made breakthrough after breakthrough in artificial intelligence (AI) over the last decade, reproducing sophisticated cognitive behaviors from advanced gameplay in board games like Go (Silver et al.,, 2016), Chess (Silver et al.,, 2018), and Diplomacy (FAIR et al.,, 2022), to generating stunning images from arbitrary natural language instructions (Betker et al.,, 2023), to catalyzing achievements in mathematics (Romera-Paredes et al.,, 2024), science (Jumper et al.,, 2021), engineering (Merchant et al.,, 2023), and medicine (Omiye et al.,, 2023; Singhal et al.,, 2023). The large language models (LLMs) powering AI products like chatGPT (Brown et al.,, 2020; Bubeck et al.,, 2023; OpenAI et al.,, 2024) have shown remarkable proficiency in cognitive domains previously regarded as uniquely human, such as natural language syntax (Linzen and Baroni,, 2021), critical reasoning and argumentation (Herbold et al.,, 2023), and computer programming (Chen et al.,, 2021; Bubeck et al.,, 2023).
In humans, these creative cognitive behaviors have traditionally been explained in part by postulating a special cognitive property: compositionality. In particular, the open-ended systematic and productive generalization behaviors exhibited by humans are often modeled by the compositionality of classical cognitive architectures, which intrinsically guarantees that they can redeploy familiar constituents in novel constructions (Fodor and Pylyshyn,, 1988). Philosophers and cognitive scientists have argued that compositionality should be considered non-negotiable in this respect, the “only game in town” for explaining the power and flexibility of human intelligence across a number of domains, particularly natural language syntax and semantics (Fodor,, 1975, 2000).
Unlike classical symbolic AI systems, however, the neural networks powering these recent advances are not designed to embody a strong assumption of compositionality, relying instead on acquiring capabilities by learning statistical regularities from data (Fodor and Pylyshyn,, 1988; Smolensky et al., 2022a, ; McClelland et al., 1986b, ; Rumelhart et al., 1986b, ). Whereas compositionality is integral to certain fundamental design features of classical computational architectures (e.g., predicate-argument structure, variable binding, constituent structure), these features are not native to standard DNNs. Traditionally, this has been lodged as an objection to the viability of neural networks as cognitive models (Fodor and Pylyshyn,, 1988). However, the advances in new DNNs’ capabilities in domains traditionally thought to require compositionality (e.g., language) calls this into question. Have today’s models overcome previously hypothesized limitations, or are they merely imitating compositional behaviors through useful but imperfect heuristics? Does the fact that DNNs reproduce sophisticated cognitive behaviors demonstrate that there is a new game in town and that compositionality is negotiable after all? Or have modern DNNs achieved genuine compositionality by non-classical means?
In this chapter, we’ll review and contextualize recent studies of compositionality in DNNs, with a particular focus on language. We begin in section 2.1 with a historical overview of the explanatory role of compositionality in philosophy, linguistics, and psychology, outlining its evolution from a property of interest primarily to logicians and philosophers to an empirically motivated psychological construct central to contemporary cognitive science. We then briefly describe the rise of connectionism in the 1980s and 90s (section 2.2) and the challenge that compositionality poses for neural networks (section 2.3). With this crucial context in place, we move to today’s deep neural networks, providing a short overview of recent advances, with an emphasis on the transformer architecture (section 3). We review key recent breakthroughs in replicating compositional behaviors in DNNs (section 4), and argue that architectural inductive biases (section 4.2) and metalearning (sections 4.3 and 4.4) offer novel approaches to reproducing the behavioral signatures of compositionality in neural networks. To close, we discuss potential implications that this work may have for our understanding of human compositionality, and suggest avenues for future empirical and philosophical work (section 5).
2 Compositionality in Context
2.1 From Frege to Fodor
Though the compositionality of natural language and thought is a central commitment of contemporary cognitive science, its current explanatory role was not always evident. Appreciation of the significance of compositionality first emerged at a time when philosophy, psychology, logic, and linguistics were still intertwined in an undifferentiated study of the mind (Pagin and Westerståhl,, 2010), and it took nearly a century of philosophical investigation to cleanly excavate the question(s) that the compositionality principle purports to answer. [^1] Compositionality can be traced even further back in the history of philosophy, to Aristotle and philosophers in the Indian tradition (Pagin and Westerståhl,, 2010), but our focus here will be on the analytic tradition, within which this notion achieved its current position of prominence and its very name (Putnam,, 1995; Katz and Fodor,, 1963).
Gottlob Frege made indispensable contributions to the disentangling of these domains (Dummett,, 1993) and was arguably the first philosopher to appreciate the explanatory potential of the compositionality principle. In contemporary terms, the compositionality principle states that the meaning of a complex expression is determined by the meaning of its parts and the way they are combined. There are a number of different ways of precisifying this principle, [^2] For discussion, see Pagin and Westerståhl, (2010) and Partee, (2004). and two importantly different ways of interpreting its significance—either as a claim about meaning itself or as a claim about the psychological processes involved in generating and grasping meanings. On the first interpretation, the compositionality principle asserts that the meaning of a complex expression is exhaustively determined by (or supervenes on) the meanings of its parts and their mode of combination. [^3] On this reading, the compositionality principle is a claim about the metaphysics of meaning, akin to the mind-brain supervenience thesis, which asserts that one’s psychological states are determined by their physical or neural states (Kim,, 1981). This first sense may be less familiar to a cognitive science and machine learning audience, but it has played a key role in long-standing debates about compositionality within philosophy. On the second, it expresses a psychological hypothesis about how language users themselves determine this meaning—namely, by assembling the meanings of the parts in accordance with tacitly known rules (Dummett,, 1996).
Though Frege clearly appreciated the significance of compositionality and anticipated key contemporary arguments in its favor (Frege,, 1956, 1980), it is not clear which of these two construals of the compositionality principle, if any, he intended. In fact, though it been dubbed “Frege’s Principle” (Cresswell,, 1973), Frege never explicitly formulated the compositionality principle himself. The first philosopher to do so was Frege’s one-time student, Rudolf Carnap, who named it after his former teacher (Carnap,, 1988). [^4] There is significant scholarly dispute over Frege’s views on the compositionality principle (Pelletier,, 2001; Janssen,, 2012). Janssen suggests that it would better be called “Carnap’s Principle.”
Carnap was the leading philosopher of the logical empiricism movement and, among other contributions, paved the way for what was arguably the central research program in 20th century formal semantics (Carnap,, 1988; Montague,, 1974). Carnap employed "Frege’s Principle" as a crucial theoretical assumption. However, like Frege before him, Carnap was deeply skeptical that these formal languages could be treated as models of natural (“vernacular”) languages (like German, English, or Mandarin) (Carnap,, 1937). His theoretical ambition was not to explain everyday language use or uncover underlying psychological mechanisms, as a contemporary theorist might expect, but to develop artificial languages ideally suited to mathematical and scientific purposes (Harris,, 2017). Compositionality was an indispensable feature of these artificial languages, but Carnap was well aware of the idealizations that they involved—abstracting away from central features of natural language, like context-sensitivity and ambiguity—and self-consciously avoided putting them forward as models of natural language. [^5] This is why Harris, (2017) frames Carnap and other ‘ideal language philosophers’ as part of the “pre-history” of natural language semantics, rather than its history proper. This held for philosophers working in the ‘ideal language tradition’ more broadly. They appreciated the power of mathematical logic and the importance of the compositionality principle, but did not advance their formal theories as psychological models or treat compositionality as an empirical hypothesis about human linguistic processing.
Ordinary Language Philosophy, the other major branch of 20th century analytic philosophy’s ‘linguistic turn’ (Rorty,, 1992), criticized the ideal language philosophers for this abstraction and focused their analyses on the messy realities of natural language use (Ryle,, 1949). Leaving aside the movement’s grander ambitions (Hacker,, 1975; Horwich,, 2013), philosophers in the ordinary language tradition made clear-cut contributions to linguistic theorizing, from J.L. Austin’s work on speech acts (Austin,, 1975) to H.P. Grice’s pioneering contributions to the field of pragmatics (Grice,, 1991). However, these philosophers were either skeptical or silent about the compositionality of natural language. To take an illustrative example, Gilbert Ryle, a leading ordinary language philosopher and philosophical behaviorist, rejected the compositionality principle and the ‘building block’ theory of meaning in which it was at home (Ryle,, 1957). Following Wittgenstein, (1973), Ryle argued that linguistic meaning is determined by use (Ryle,, 1957, 1964), [^6] This Wittgensteinian perspective on linguistic meaning inspired J.R. Firth’s distributional semantics (Firth,, 1957), a key historical antecedent of the statistical approaches powering today’s large language models (Brunila and LaViolette,, 2022). and that use is highly contextual—trying to extract a single, unified contribution that a given word makes to each of the sentences (or speech acts) in which it may occur would be a methodological dead-end. Put simply, in Ryle’s view, the use theory of meaning runs counter to compositionality, so the compositionality principle should be rejected.
The tension between these approaches to the study of language persisted into the early 1960s, with ideal language philosophers developing compositional models, but abstracting away from natural language use, and ordinary language philosophers advancing analyses of natural language, but rejecting or simply ignoring the principle of compositionality. This contrast formed the philosophical backdrop for Noam Chomsky’s revolution in the field of syntax (Chomsky,, 1957, 1965), which would fundamentally alter the role of compositionality in linguistic theorizing.
Chomsky rejected the key assumption that both groups of philosophers shared—that formal mathematical theories could not be treated as models of ordinary language. This allowed him to synthesize their approaches, harnessing the “methodological toolkit” developed by philosophers interested in formal logic, from Frege on (Harris,, 2017), but taking them seriously as models of the syntax of natural languages (Chomsky,, 1957; Katz,, 1971). Though Chomsky did not focus on questions of meaning, and would eventually express some skepticism regarding the compositionality of natural languages (Chomsky,, 1975), the synthesis that his work enacted was a crucial step in the emergence of the compositionality principle as an empirical hypothesis about human linguistic processing. Indeed, it was promptly extended to formal semantics by Richard Montague and other key contributors (Katz and Fodor,, 1963; Davidson,, 1965). Montague explicitly argued that there is “no important theoretical difference between natural languages and the artificial languages of logicians,” allowing him to treat rule-based semantic systems of the kind pioneered by Carnap as models of human linguistic competence (Montague,, 1974, p. 15). [^7] Montague signals his alignment with Chomsky on this key point: “I consider it possible to comprehend the syntax and semantics of both kinds of languages within a single natural and mathematically precise theory. On this point I differ from a number of philosophers, but agree, I believe, with Chomsky and his associates” (Montague,, 1974, p.15).
It would be hard to overstate the impact that this shift had on linguistic theorizing. In effect, it transformed formal syntactic and semantic theories from abstract objects of study for logicians, mathematicians, and philosophers into psychological constructs that underlie and explain the linguistic capacities of even the youngest children capable of speech.
With this shift, the compositionality principle took on a newfound significance and key role in explaining human linguistic behavior. Chomsky drew attention to the creativity of human language use and its open-ended, potentially infinite expressive capacity, which must be achieved through finite means (Chomsky,, 1966, 2020; Frege,, 1980). To use a classic example discussed by Katz and Fodor, (1963), even in the most routine linguistic interactions we regularly encounter novel sentences, yet fluidly and effortlessly comprehend their meaning and generate appropriate responses. This open-ended creative capacity seems to cry out for explanation. [^8] Chomsky suggests that some 17th century philosophers, like Descartes and Wilhelm von Humboldt, recognized this explanatory challenge and anticipated some of his own ideas (Chomsky,, 1966). It is also prefigured by Frege, who wrote in 1914 that, “the possibility of our understanding sentences which we have never heard before rests evidently on this, that we can construct the sense of a sentence out of parts that correspond to words… Without this, language in the proper sense would be impossible… we would always be restricted to a very narrow area and could not form a completely new proposition” (Frege,, 1980). How are humans capable of understanding sentences that they have never before encountered?
The answer, in one word, is compositionality. If the meaning of an unfamiliar sentence is determined by the meaning of its parts, their mode of combination, and nothing else, then knowing the meanings of component words and the grammar of the language will suffice for grasping this novel meaning. Appealing to tacit knowledge (Fodor, 1968a, ; Evans,, 1985) thus provides an explanation of the human capacity to understand novel sentences—we derive the meaning of novel constructions by recombining known components according to known syntactic rules.
In fact, it can seem that tacit knowledge of a compositional grammar is the only adequate explanation of this capability. Chomsky argued that humans’ creative capacity is essentially unbounded (Chomsky,, 1957). To take a classic example, if we idealize away from constraints of memory and attention (as well as the finitude of a human lifetime!), a compositional subject could go on adding new conjuncts to the end of a well-understood sentence ad infinitum. Provided that they understand the meaning of the new conjunct (and conjunction itself), they should also be able to understand these newly formed sentences.
It is this open-endedness that makes an internalized compositional system seem like “the only game in town.” Alternative explanatory approaches of the time, like behaviorism, struggled to account for this form of unbounded extension. A behaviorist account of such capacities in terms of simpler psychological mechanisms like stimulus-response association (via classical or operant conditioning), would fail to explain these generalization behaviors involving previously unencountered stimuli. For Chomsky, Montague, Fodor, and others, the clear solution was to posit further internal compositional mechanisms that support these behavioral capacities.
This abductive inference to the best explanation forges a close link between creativity, productivity, and compositionality. Particular observations about productive or systematic behaviors in humans motivate the ascription of internal compositional mechanisms (which behaviorists would never have countenanced) to ordinary speakers. Such arguments commit theorists to interpreting the principle of compositionality as a psychological, rather than a metaphysical, thesis. This point deserves emphasis. To serve as a robust explanation of human language comprehension and production, this compositional system must actually get deployed by humans when they generate the behaviors that the compositionality principle purports to explain. [^9] This observation may seem obvious, but it is in fact easy to overlook. Even Davidson, one of the earliest and most influential proponents of compositional natural language semantics, fails to fully appreciate the point. For a detailed discussion of this tension in Davidson’s work, see Heck, (2013). For the proposed explanation to work, it is not enough that the language itself admits of a compositional analysis—after all, one and the same language may admit both compositional and non-compositional models.
To see the point, consider a simplified ‘toy’ language (Evans,, 1985) with only 10 names (a, b, c…) and 10 predicates (F, G, H…), generating 100 possible sentences (Fa, Fb, Ga…). This language can be provided with a compositional model, along familiar lines. However, every admissible sentence in this language could also be generated by a very different, though ‘extensionally equivalent,’ model that has 100 different axioms specifying the meaning of each. How can we tell which of these models a speaker of this minimal language possesses? Many philosophers, from W. V. O. Quine, (1970) to Crispin Wright, (1981), were skeptical that such questions could be made empirically or behaviorally tractable. Evans, (1985), however, argues that there are clear empirical considerations that justify the ascription of the compositional system rather than the non-compositional one. In principle, we could scrutinize the speaker’s internal states and attempt to detect compositional structure. In practice—and in keeping with the methodological assumptions of researchers in the Chomskyan tradition—we can instead rely on a wider class of behavior—compositional generalization behavior involving the introduction of novel names or predicates (if you introduce a new sentence, Fz, will the speaker be able to understand Gz and Hz?). Evans’ proposal about how to operationalize compositionality is very similar to those used in contemporary machine learning research (see section 4.1), and it demonstrates the central role that linguistic behavior plays in motivating cognitive explanation in terms of the compositionality principle.
These theoretical developments close the gap between early philosophical theorizing about compositionality and the ongoing debates in contemporary cognitive science. Compositionality is understood as an explanatory psychological property, operationalized in terms of the specific kinds of productive and systematic generalization behaviors that humans exhibit (Evans,, 1985). In each case, the characteristic ‘behavioral signatures’ of compositionality are taken as evidence diagnostic of internalized compositional mechanisms, which serve to explain the behaviors in question. As Chomsky often emphasizes, this means that even the most abstract reaches of formal linguistics are properly understood as a branch of cognitive psychology (and ultimately biology, though developing an account of how symbolic systems are implemented in the brain is sometimes treated as a subsidiary enterprise; Chomsky,, 1968).
Two further consequences of the Chomskyan revolution deserve mention before concluding this historical overview. First, the rise of this new explanatory paradigm brought with it a notable resurgence of nativism, as Chomsky and others argued that the linguistic data to which children are exposed is insufficient for recovering the kind of fine-grained syntactic structures necessary for explaining human linguistic competences (Laurence and Margolis,, 2001). As we will briefly discuss in 5.2, the rise of DNNs and LLMs may make novel contributions to this long-standing philosophical issue (Buckner,, 2018, 2019; Buckner, 2023b, ). And second, though we have framed the discussion in terms of the meaning of sentences throughout this section, this period saw a distinct shift in focus from language to thought. This widely discussed ‘cognitive turn’ (Miller,, 2003) overthrew behaviorist strictures on psychological explanation (Fodor, 1968b, ) and transposed many of the debates central to the linguistic turn into questions about the structure of cognition itself (Burge,, 2007). Compositionality was no exception. Jerry Fodor influentially argued that the compositional structure of natural languages is a reflection of the compositional structure of thought, reintroducing the language of thought hypothesis to the philosophical mainstream (Fodor,, 1975; Pelletier and Roques,, 2017). On this view, human cognition is compositional from the ground up—it is a feature of the mind’s basic operating system, not just an artifact of one specific domain (language).
Across this historical evolution, compositionality arrives in the position familiar to cognitive scientists today: an empirically motivated, non-negotiable property of both language and thought and the best explanation of the creativity, systematicity and productivity of human cognition. Throughout his career, Fodor would wield compositionality like a sledgehammer to smash rival theories of meaning and defend the language of thought hypothesis. In fact, he essentially runs the contrapositive of Ryle’s argument above—because the compositionality principle is non-negotiable and the use theory of meaning does not ensure it, the use theory of meaning must be false (Fodor and Lepore,, 2001; Fodor,, 2003). Over the years, he employed the same logic to attack conceptual role semantics (Fodor and LePore,, 1993), prototype theory (Fodor and Lepore,, 1996), and most importantly for our present purposes, to neural network models of cognition (Fodor and Pylyshyn,, 1988).
2.2 The Rise of Connectionism
In the 80s and 90s, a series of key advances in neural networks, such as the invention of Hopfield networks (Hopfield,, 1982) and the backpropagation algorithm (Rumelhart et al., 1986a, ), led many to abandon this classical, Fodorian picture and turn to an alternative view of the mind (McClelland et al., 1986b, ; Rumelhart et al., 1986b, ). The connectionist movement was initially motivated by considerations other than compositionality, such as the basic computational structure of biological brains. Proponents sought to demonstrate that sophisticated cognitive behaviors could emerge in networks of simple interconnected units (or “neurons”). Though these units abstract away from a great deal of what is known about biological neurons, they implement a similar kind of computation: each unit performs a weighted sum of its inputs and applies a threshold or nonlinearity, much as biological neurons integrate dendritic inputs and spike at a certain threshold (Hodgkin and Huxley,, 1952; McCulloch and Pitts,, 1943; O’Reilly et al.,, 2012). Early work demonstrated how these networks could learn by incrementally updating their weights (or “synapses”) to improve overall performance on a given task (Rumelhart et al., 1986b, ).
Connectionists sought to construct models that explained behavioral phenomena without building in classical representational elements (e.g., rules, variables, symbols, propositions, constituent structure), instead emphasizing how useful representations could emerge in a neural networks’ distributed patterns of activity over the course of learning (Rumelhart and McClelland,, 1986). This approach stood in tension with existing explanations based on the principle of compositionality, since standard distributed representations have no natural notion of “parts” or “constituents” (although key later work by Smolensky,, 1990, would explore this possibility). The network’s objective is to discover and deploy whatever representational elements are most useful for solving a given problem. These emergent representations may not align with the kinds of features a linguistic or cognitive theorist might put forward, and may not even afford succinct interpretations in psychological terms at all (Smolensky,, 1986). In fact, as we have discussed elsewhere (McGrath et al., 2023a, ), many early connectionists were eliminativists about classical symbolic constructs (Rumelhart and McClelland,, 1986; Ramsey et al.,, 1990; McClelland and Patterson,, 2002), arguing that the rise of neural networks repudiated traditional approaches to cognitive science and constituted a paradigm shift for the field (McClelland et al., 1986a, ).
Neural networks thus seemed to offer a radically new approach to the study of cognition, providing novel explanations for many different kinds of behavioral phenomena (Rumelhart et al., 1986b, ), including perceptual processes (McClelland and Rumelhart,, 1981), memory processes (Hopfield,, 1982; McClelland et al.,, 1995), executive and decision-making processes (Cohen et al.,, 1990; Miller and Cohen,, 2001), and language processes (Dell,, 1986; Rumelhart and McClelland,, 1986). The emergence of the connectionist movement generated a great deal of excitement and controversy across the cognitive sciences, including within philosophy (Ramsey et al.,, 1991), where Fodor was waiting (with his sledgehammer).
2.3 The Compositionality Challenge
In response to the threat posed by connectionism, Fodor penned a long critique of the movement with the psychologist, Zenon Pylyshyn (Fodor and Pylyshyn,, 1988). Their highly influential article, “Connectionism and Cognitive Architecture: A Critical Analysis,” develops a dilemma for proponents of neural network models of cognition. [^10] Framing Fodor and Pylyshyn’s argument as a dilemma is common in the literature (Millière and Buckner, 2024a, ) and is consonant with the earliest connectionist replies to their article (Smolensky,, 1991), but one could argue that the two horns of this dilemma really comprise a single objection—that neural networks cannot explain compositionality. Our goal here is not to engage in ‘Fodorology.’ Dividing what we will call the ‘compositionality challenge’ into two distinct horns is, in our view, a useful way of distinguishing two importantly different objections to DNNs, which often get misleadingly run together. The first horn of this dilemma is the claim that core properties of neural networks make them incapable of replicating the behavioral signatures of compositionality, rendering them empirically inadequate as models of human cognition. This claim has real predictive upshot—if it is correct, neural networks should fail to match human performance on the specific tasks that demand compositional generalization (Marcus,, 1998). The second horn, which we will call the ‘mere implementation’ objection, is subtler and involves more overtly philosophical considerations. It alleges that, even if DNNs were to replicate the behavioral signatures of compositionality, this would only mean that they had merely implemented a classical symbolic architecture (McGrath and Russin,, 2024), rather than providing a genuine alternative explanation. Taken together, these two horns comprise the compositionality challenge for artificial neural networks.
This first horn of Fodor and Pylyshyn’s argument draws on two key cognitive properties that are taken to require compositionality: productivity and systematicity. As we have seen, classical symbolic accounts provide a natural mechanism for productively generating an unbounded set of novel thoughts through the recombination of a finite pool of primitive elements. Connectionist architectures, on the other hand, do not seem to build in the kind of combinatorial constituent structure capable of supporting this form of unbounded productivity. They also struggle to capture the systematicity of cognition—the apparently intrinsic connection between the ability to comprehend and generate specific sets of thoughts/sentences. To use the classic example, it seems that anyone capable of understanding the sentence “John loves Mary” should be capable of understanding the sentence “Mary loves John.” The classical account again provides a natural explanation—the meaning of the second sentence is a function of the very same components and mode of combination as the first. Therefore, any compositional learner capable of understanding the first will be able to redeploy this knowledge in understanding the second. [^11] This systematicity challenge differs from the challenge provided by productivity. For one thing, it does not rely on idealizing to an infinite cognitive capacity, and so may seem to yield the same conclusion from a weaker premise (a key motivation for introducing the argument in the first place—see Fodor,, 1987, p. 147-154). For important qualifications, however, see Szabó, (2022). Fodor and Pylyshyn argue that in a connectionist network without constituent structure, on the other hand, each of these individual acts of understanding would be an independent capacity. There would therefore be no inherent reason why they should hang together—a network trained on one could easily misunderstand the other. For both productivity and systematicity, Fodor and Pylyshyn’s conclusions are the same—without underlying compositional mechanisms, connectionist networks will fail to replicate these core features of human language and cognition.
The second horn of Fodor and Pylyshyn’s dilemma draws on the possibility of implementing a classical symbolic system in an artificial neural network. Even if a connectionist network did manage to exhibit systematic and productive behaviors, they argue, this would have to be because it had merely implemented a classical symbolic system, undermining any claim to revolutionary implications or novel insight into the nature of the mind (for further discussion, see section 5.3). Thus, the connectionist is left with no way out: either their models are empirically inadequate, or they are mere implementations, which may prove informative for theorists interested in how symbolic systems are realized in the human brain (McLaughlin,, 1993), but will not contribute to higher-level cognitive theory.
As should be clear, this argument is a species of the genus described in section 2.1—compositionality is treated as an empirically motivated psychological construct, necessary for the explanation of certain key human behaviors.
In the context of connectionism, however, the compositionality challenge ignited a heated debate and drew a wide array of rebuttals and clarifications (van Gelder,, 1990; Matthews,, 1994; Ramsey et al.,, 1990; Horgan and Tienson,, 1991). Some connectionist researchers—most notably Paul Smolensky—tried to meet this ‘mere implementation’ objection head on by building neural networks with internal compositional structure that did not merely recapitulate classical architectures (Smolensky,, 1990, 1991). This led to further debate about just what kind of explanation compositionality offered and what internal compositional structure was rightly considered ‘non-negotiable’ (MacDonald and MacDonald,, 1991; van Gelder,, 1990). Robert Hadley, (1997) observed that the demand that systematicity be a matter of nomic necessity, an inherent or intrinsic element of the architecture, is too strong, since even a classical symbolic architecture could be programmed to exhibit failures of systematicity. And David Chalmers, (1993) pointed out that Fodor and Pylyshyn’s argument proves too much and so cannot be correct, as formulated. As all parties to the debate acknowledge, connectionist networks may implement classical symbolic architectures. However, Chalmers noted, Fodor and Pylyshyn’s argument for the first horn of the dilemma hinges on basic features of neural networks that are present even in such connectionist implementations of classical architectures. Yet nowhere do they specify an ‘escape clause’ by which these implementations are exempt from the first horn of the argument (Chalmers,, 1993). Thus, if their initial argument were correct, it would prove that even classical symbolic architectures (implemented in neural networks) lack productivity and systematicity, a conclusion that is clearly untenable and self-defeating.
Our primary goal in this paper—and the sole focus of the review in section 4—is to survey recent findings suggesting that contemporary DNNs have plausibly surmounted the first horn of this dilemma. In 5.3, we also briefly evaluate whether this means that they fall on the second horn (for further discussion, see McGrath et al., 2023a, ; McGrath and Russin,, 2024).
Over time, the debates over connectionism waned, moving from the center to the periphery of cognitive science. Neural networks, however, continued to progress, and exploded into new prominence with the inception of the deep learning revolution (Krizhevsky et al.,, 2012; LeCun et al.,, 2015; Goodfellow et al.,, 2016), instilling the debates over connectionism and compositionality with renewed importance.
3 Neural Networks in the Age of Deep Learning
The primary factor that differentiates modern neural networks from their connectionist predecessors is their scale: modern networks are deeper (i.e. more layers), have many orders of magnitude more learnable parameters (weights), and are trained on many orders of magnitude more data. This is possible due to a combination of advances in hardware (e.g., GPUs) and the availability of large amounts of data from the internet (e.g., Wikipedia). These are merely quantitative differences, but they have led to qualitative increases in their capabilities (Brown et al.,, 2020; Caballero et al.,, 2023; Kaplan et al.,, 2020; Wei et al., 2022b, ).
While scale has perhaps been the single most important factor, there have also been advances in areas such as network architecture, where the most notable development has been the introduction of the transformer (Vaswani et al.,, 2017). In this section, we will briefly review transformers. [^12] Readers already familiar with transformers can skip directly to section 4. Those interested in learning even more about transformers may want to read the original paper in which they were introduced (Vaswani et al.,, 2017), as well as helpful introductory texts (e.g., Millière and Buckner, 2024a, ), and some of the many helpful blog posts available online (e.g., Alammar,, 2018).
Invented in 2017, the transformer has become the dominant architecture in natural language processing models (e.g., in the LLMs powering chatGPT) and has also become common in vision models (Dosovitskiy et al.,, 2021; Dehghani et al.,, 2023), multimodal models (Xu et al.,, 2023), and even reinforcement learning models (Li et al.,, 2023). Transformers take a sequence of vectors as inputs and produce a sequence of vectors as outputs. In the case of language models, each input vector will be an embedding of an individual word (or to be more precise, a sub-word token, which can be a whole word or a part of a word; Sennrich et al.,, 2016; Radford et al.,, 2019). One of the main advantages of the transformer architecture is that it processes input sequences (e.g., sentences, paragraphs, or documents of text) in parallel, avoiding the computational cost of recurrent neural networks (RNNs), which need to compute intermediate states serially across the sequence (see Figure 1A). RNNs maintain and update a single hidden state as the input sequence is processed serially, effectively requiring them to continuously compress the entire past (i.e., anything from prior inputs that could be relevant to future outputs) into this single hidden state. Transformers have no such constraint, as encodings at each position in the sequence can be updated by information in arbitrary other positions via attention (see Figure 1B).
Figure 1: Transformer basics. A) Comparison with recurrent neural networks (RNNs) in the language modeling setting (predicting the next word). A fundamental difference between RNNs and transformers is that RNNs must compress the entire past into a single hidden vector. To see this, notice how the only way information from the word ‘Attention’ can get to the position of the word ‘you’ is via the hidden states at the intermediary positions (highlighted with red arrows). In the transformer, information from the word ‘Attention’ can get to the position of the word ‘you’ directly at any layer via the attention mechanism (single red arrow). Attention weights are depicted by the thickness/darkness of the lines (which have been chosen arbitrarily for illustration purposes). B) Attention mechanism. At each position in the sequence, queries (blue), keys (green) and values (purple) are determined by multiplying inputs by learnable weight matrices ($W_{q}$ , $W_{k}$ , $W_{v}$). For simplicity, key and value are not shown at position 4, and queries are not shown at other positions. Queries (blue) are compared to keys (green), and the result determines the attention weight. A good match means the value vector (purple) at that key’s position will receive a large weight (shown by darker lines).
The attention mechanism (Bahdanau et al.,, 2014; Vaswani et al.,, 2017) is so named because it allows representations at a given position in the sequence to selectively filter (“attend to”) the information it receives from other positions (like how humans selectively filter information in the environment in order to accomplish a goal; Lindsay,, 2020). In self-attention, a layer of encodings of a sequence can attend to the previous layer of encodings of the same sequence. This is opposed to cross-attention, where a decoder attends to a given sequence while producing a new sequence (e.g., in another language during a translation task). Each layer of the transformer has a number of heads, which process the input vector at each position in the sequence, multiplying it by three separate learnable matrices to form a query, a key, and a value vector. Intuitively, the query vectors can be thought of as soliciting information that is needed at a given position (e.g., what is this verb’s direct object?), the key vectors as advertising what information is available at a given position (e.g., that this noun might be a direct object), and the value vectors as representing the content that will actually get transmitted if there is a good match between key and query (e.g., that this direct object is “John”). Thus, the keys and the queries fully determine how information will be transmitted across positions in the sequence according to how well they match, as measured by their dot product:
$\text{Attention}(Q,K,V)=\text{softmax}(\frac{QK^{T}}{\sqrt{d}})V$ (1)
Here, the matrix multiplication $QK^{T}$ computes a matrix where each element is the dot product of a particular query-key pair (normalized by the square root of the dimension of the vectors, $d$), and the softmax ensures that each row of the matrix will be a probability distribution, which determines the weights that multiply the value vectors in $V$. The final output at a given position of one attention head, then, is an attention-weighted average of the value vectors from all positions, where the attention weights are determined by how good the match is between that position’s query and another position’s key. [^13] By default, the “residual stream” (Elhage et al.,, 2021) of the transformer does not mix information across positions; for information to get from position i to position j, the key vector at position i must have a large dot product with the query vector at position j. This will give a large value for the input to the softmax at entry (i,j) which will mean a higher proportion of the final output at position j will be constituted by the value vector from position i. In effect, the attention mechanism allows the network to combine and integrate information across the sequence in arbitrarily flexible yet selective ways.
Importantly, the designer of a transformer doesn’t actually specify in advance how the attention should operate (i.e., what kinds of information will go into the queries, keys, or values for a given head)—these will be resolved in the course of the optimization process, where the network will discover the best ways of using the learnable parameters in the matrices of the attention mechanism to succeed on its task (i.e., to reduce its loss). Some recent work has attempted to understand what kinds of representations and attentional strategies these networks end up discovering. For example, some attention heads in language models seem to be sensitive to syntactic structures (Manning et al.,, 2020). Other work has found “induction heads” (Olsson et al.,, 2022) that search the context for previous occurrences of the current token and encode what came after it, thus allowing the network to “reason by induction” to predict that the same pattern will repeat again. Rapid progress has been made in methods for discovering the circuits underlying a given capacity in trained LLMs (“mechanistic interpretability”; Conmy et al.,, 2023; Elhage et al.,, 2021; Merullo et al.,, 2024; Olsson et al.,, 2022; Wu et al.,, 2023), but it is an important open question whether the mechanisms learned by LLMs in the course of training will all afford these kinds of succinct algorithmic descriptions (Arrieta et al.,, 2019; Crook et al.,, 2023; Lipton,, 2017).
Thus, the fundamental spirit of connectionism is alive and well in the deep learning revolution: practitioners don’t build into their networks the representations and processes that they will use to solve a given task, instead allowing useful representations to emerge in their distributed patterns of activity learned through trial and error on data. In some ways, modern deep learning researchers have taken this idea to new extremes. Whereas older work in machine learning used highly engineered features to make the learning problem easier for a downstream learning system such as a shallow neural network, modern practitioners skip the feature-engineering step entirely and train deep networks directly on raw images or text, allowing the early layers to discover whatever features will be most useful for downstream layers through “end-to-end” learning (Goodfellow et al.,, 2016). This “bitter lesson” (Sutton,, 2019)—that in high-dimensional domains like vision or language, deep neural networks will usually find better solutions than the ones that even the cleverest engineers could come up with—stands in direct conflict with the spirit of Chomskyan and Fodorian nativism, which emphasizes that certain properties (e.g., compositionality) must be necessary consequences of the specific kinds of representations that are built into the architecture (Fodor and Pylyshyn,, 1988; Fodor,, 1975).
4 Review
Although these recent developments in DNNs have led to remarkable performance gains across many domains, it is unclear a priori whether any of the specific advances (e.g., increasing scale, the transformer architecture) should have a significant impact on their underlying compositionality. The features of neural networks that led theorists like Fodor and Pylyshyn, (1988) to conclude that they were not compositional seem prima facie to be largely unchanged in modern deep networks: transformers are still extremely unbiased learners with no explicit rule-like representations or built-in sensitivity to constituent structure. [^14] While most agree that the transformer is a very unbiased architecture (e.g., it is in some ways even more unbiased than the convolutional neural network; Dosovitskiy et al.,, 2021), some have argued that the attention mechanism might play a special role in the compositional behaviors observed in transformers (Smolensky et al., 2022a, ). As we have seen, accounting for compositionality is a pressing theoretical challenge—if neural networks are to provide illuminating accounts of human cognitive capacities, they must be able to model the creative, productive, and systematic generalization behaviors exhibited by humans.
Our central question in this review will be an empirical one: can contemporary neural networks replicate the behavioral signatures of compositionality? We will argue that although many current neural networks (including ones leveraging modern architectures such as the transformer) still seem incapable of capturing these behavioral signatures, a few novel approaches have shown that it is possible to endow neural networks with the properties that are required to produce them. These results cast doubt on the empirical predictions made by classical theorists that neural networks would never be capable of generalizing systematically or productively, and suggest that DNNs may meet the challenge posed in the first horn of Fodor and Pylyshyn’s (1988) dilemma.
4.1 Operationalizing Compositionality
An empirical investigation of compositionality in neural networks requires that we operationalize it by specifying exactly what kinds of behaviors it is supposed to afford—we need a set of behavioral measures that allow us to determine whether a given learner is compositional (Garson,, 1994). Broadly speaking, compositionality has been evaluated in DNNs by measuring the degree to which they generalize under specific conditions. This reflects the explanatory role played by compositionality for classical theorists (Katz and Fodor,, 1963; Evans,, 1985): compositionality is a property of the underlying cognitive system whose existence we can infer precisely because it explains certain kinds of generalizations (e.g., systematic and productive ones) that we observe in the system’s behavior.
One might think that in order to determine whether a given learner is really compositional, its internal representations and/or processes need to be scrutinized in addition to its behavior. For example, some work has attempted to measure the extent to which a model’s internal representations capture ground-truth compositional or constituent structure (Andreas,, 2019; Lepori et al.,, 2023; Lewis et al.,, 2023; McCoy et al.,, 2019; Russin et al., 2021a, ; Soulos et al.,, 2020; Todd et al.,, 2024). We think this work is interesting and briefly discuss it in the final section (5.3). However, we see this interpretive work as being relevant not to the first horn of the compositionality challenge, but to the second horn—the question of whether a system is producing compositional behaviors by ‘merely implementing’ a classical architecture. As we discussed in section 2.3, the first horn of the dilemma, in our view, is an empirical question that can be answered with reference to behavioral evidence, while the second horn of the dilemma requires both empirical investigation into the inner workings of a given neural network and philosophical work to determine what counts as a ‘classical architecture’ (Garson,, 1994), whether a given network is implementing one, and what should count as a satisfactory explanation of systematic and productive behavior.
The behavioral tasks used to evaluate the compositionality of DNNs typically operationalize it as the ability to generalize from a given set of training samples to a given set of test samples. These train and test sets are crafted such that a compositional learner ought to be able to infer the underlying grammatical structures or rules that govern the training examples, and then leverage that knowledge to systematically generalize to the test samples. [^15] An alternative way to operationalize compositionality is to test sample efficiency, or the number of training samples it takes to achieve a certain level of test performance (Bahdanau et al.,, 2018). This makes a similar assumption that a compositional learner should be able to infer the rules governing the data from relatively few samples. For example, the model might be trained on the meanings of instructions using novel verbs, like “dax,” “flug,” and “flug twice,” and then tested on an unseen instruction “dax twice.” If the learner is compositional, it should be able to extract the meanings of “dax,” “flug,” and “twice” from the training samples, as well as knowledge of how adverbs modify verbs in the underlying grammar. [^16] It would probably require more than the three examples given above to truly infer their underlying structure, but we have chosen to keep it brief for illustrative purposes. If the model infers this underlying structure correctly, it should be able to generalize to the test sample (“dax twice”) by composing the familiar meaning of the verb “dax” with the familiar meaning of the adverb “twice” according to the known grammatical rules (Lake and Baroni,, 2018). On the other hand, a non-compositional learner—for example, a stimulus-response learner that simply stored each training sample in a lookup table (see Figure 2)—would have no basis upon which to generalize to the novel expression “dax twice.”
Figure 2: Are neural networks stimulus-response learners, or do they learn to solve problems with compositional algorithms? Two extreme possibilities are depicted for natural language (top) and simple arithmetic (bottom). In either case, a stimulus-response learner (left) would simply memorize a lookup table where entries corresponded to entire inputs/outputs, ignoring any compositional structure. A compositional learner (right) would solve the same problems by composing familiar elements according to known rules, allowing generalization. Figure reproduced from Russin et al., 2021a .
In a seminal study on compositionality in modern DNNs, Lake and Baroni, (2018) leveraged these principles to develop the SCAN task, a sequence-to-sequence task where the model takes instructions as inputs (e.g., “walk left”), and must return the corresponding sequence of actions (e.g., TURN_LEFT WALK). The instructions were generated from a simple, finite phrase-structure grammar (about 20,000 total possible instruction sequences), and corresponding action sequences were generated by a deterministic interpretation function. To test compositional generalization, the researchers created a systematic difference between the train and test sets: one of the four primitive verbs (“jump”) was left out of the training set except in its simplest form (“jump” → JUMP). [^17] The dataset also included other generalization tests, such as a split designed to test generalization to longer sequences (Lake and Baroni,, 2018). Here we focus on the ‘add-jump’ split, which was the main compositional generalization test on which standard networks were shown to fail. This created a scenario like the one discussed above, where a model could in principle infer the underlying grammar from all of the complicated constructions containing the other primitive verbs (e.g., “walk twice,” “look twice,” etc.), and combine this with knowledge of the action associated with “jump” (JUMP) to generalize to the analogous constructions containing “jump” (e.g., “jump twice”).
The results showed that even though standard modern DNNs performed nearly perfectly when the train and test sets were randomly split (i.e., when the train and test data were randomly sampled from all possible instructions), these same networks performed quite poorly on this compositional generalization test (about 1.2% accuracy reported for the RNNs tested in the original study). This was true for all of the models they tested, and follow-up work showed that other standard architectures such as convolutional neural networks (Dessì and Baroni,, 2019) and transformers (Keysers et al.,, 2020) exhibit similar reductions in performance on the compositional test (though less extreme in the case of the convolutional neural network). Analogous experiments in humans using a modified version of the same task have shown that humans perform quite well (Lake et al.,, 2019; Lake and Baroni,, 2023). A number of subsequent machine learning studies followed up on this work, creating similar tests of compositional generalization and showing that these standard DNNs fail to generalize compositionally in domains such as semantic parsing (Kim and Linzen,, 2020), string manipulation with a probabilistic context-free grammar (Hupkes et al.,, 2020), mathematics (Saxton et al.,, 2019), natural-language translation (Dankers et al.,, 2022), and grounded language understanding (Ruis et al.,, 2020), among others.
Taken together, the studies that have evaluated the compositionality of standard DNN architectures show that while they excel at generalizing when train and test data are split randomly, they perform remarkably worse when there is a systematic difference between train and test. One way of understanding these results is in terms of classical cognitive theory—as vindicating theorists who held that compositionality should be a necessary consequence of cognitive architecture and consequently predicted that neural networks will inevitably fail to replicate the systematicity and productivity apparent in human performance (Fodor and Pylyshyn,, 1988; Marcus,, 1998).
Figure 3: Two different kinds of generalization in one dimension (top) and in higher dimensions, as in the SCAN task (bottom). When train data and test data come from the same domain, generalization to the test data only requires interpolation (top left). When test data is sampled from a different domain, generalization requires extrapolation (top right). In the bottom row, each shape represents the distribution of all possible instruction sequences in the SCAN grammar (Lake and Baroni,, 2018), which can be split into train and test in different ways. When train and test data are sampled from the same distribution, i.e., independent and identically distributed (i.i.d.), neural networks generalize well (bottom left). However, in compositional generalization tests there are systematic differences between train and test data, thus requiring out-of-domain (o.o.d.) generalization. For example, in the compositional generalization test in the SCAN task (Lake and Baroni,, 2018), the instructions containing “jump” are systematically left out of the training set (bottom right). In reality these samples would be difficult to visualize in two dimensions, but for the sake of illustration, all such instructions are depicted here as falling in the upper left quadrant of the distribution to emphasize the extrapolation required in the o.o.d. case. Additionally, each quadrant is depicted as sharing an underlying symmetry with the other quadrants, corresponding to the symmetry with the set of instructions containing the other three primitive verbs. Figure adapted from (Russin et al., 2020a, ).
Another (perhaps compatible) way of understanding these results is to take a statistical learning theory perspective (see Figure 3). Neural networks are universal function approximators (Cybenkot,, 1989; Hornik,, 1991) that assume that the train data and test data are sampled from the same distribution, i.e., that the train and test data will be independent and identically distributed (i.i.d.). Viewed through this lens, the generalization behaviors that are taken to be diagnostic of compositionality (Evans,, 1985; Fodor and Pylyshyn,, 1988) are out-of-domain (o.o.d.) generalizations (see Figure 3), and require extrapolation, rather than interpolation (Marcus,, 1998, 2018). [^18] Out-of-domain generalization is also sometimes referred to as “out-of-distribution” or “out-of-sample” generalization, and is required in any “non-stationary” environment where the data-generating distribution changes over time. Also, see Balestriero et al., (2021) for problems with this usage of the terms ”interpolation” and ”extrapolation” in high-dimensional settings. For an unbiased learner like a standard neural network, this kind of generalization should be unlikely or impossible.
The extent to which a learner is expected to generalize beyond its training distribution is the extent to which it must make additional assumptions about the world. These assumptions, called ‘inductive biases’ in statistical learning parlance, can take many forms in neural networks, including architectural inductive biases (e.g., convolutional neural networks assume translation invariance/equivariance), extra regularization terms in the loss function (e.g., weight decay, a regularization term that encourages a network’s weights to be small), etc. Standard neural networks do not make very many of these assumptions and thus tend to struggle when even the most superficial distributional shifts occur between train and test (Lake et al.,, 2017). For example, standard neural networks used in deep reinforcement learning agents trained on Atari games fail to generalize when superficial or trivial features of the game are changed (Kansky et al.,, 2017). It is not surprising, then, that they also fail in compositional generalization settings where there can be extreme distributional shifts from train to test.
However, recent work has investigated methods that endow neural networks with inductive biases that would facilitate their adaptation to these shifts and allow them to replicate the behavioral signatures of compositionality. In the following, we’ll review recent work from three different areas suggesting that neural networks can acquire such inductive biases:
  1. 1.
    Architectural inductive biases: Some work has focused on designing specialized network architectures that would facilitate the emergence of compositional generalization behavior. This can include neuro-symbolic hybrids, which combine explicit symbolic processes with neural networks, or specialized neural architectures that do not build in such explicit symbolic processes.
  2. 2.
    Metalearning: In the metalearning setting (Lake and Baroni,, 2023), an inductive bias is imparted to a neural network by training it on a specific distribution of tasks (e.g., compositional generalization tasks), endowing it with an ability to learn new tasks in specific ways.
  3. 3.
    Large-scale pretraining: When neural networks are pretrained on a large amount of data, they can acquire emergent capabilities (Brown et al.,, 2020; Bubeck et al.,, 2023). Similar to the metalearning setting, pretraining can be thought of as a way of imparting inductive biases to a learner.
In the following, we will review key findings from recent studies exploring compositionality using each of these approaches, highlighting how each challenges the traditional view of compositionality in neural networks. Our review will largely focus on language, as this has been the domain in which compositionality has traditionally played a central explanatory role (see section 2.1). However, there has also been a recent surge of interest in compositionality in vision (Lepori et al.,, 2023; Zhou et al.,, 2024), multi-modal models (Castro et al.,, 2024; Hsieh et al.,, 2023; Lewis et al.,, 2023; Ma et al., 2023b, ; Ma et al., 2023a, ; Mitra et al.,, 2023; Ossowski et al.,, 2024; Shukor et al.,, 2023), and in reinforcement learning (Bakirtzis et al.,, 2022; Jothimurugan et al.,, 2021; Liu and Frank,, 2022; Žikelić et al.,, 2023), where similar issues have been explored. Furthermore, while we have chosen to focus on the above three topics to organize our review, it should be noted that much of the recent work does not fall cleanly into any of these categories. For example, some work has investigated the effects of standard architectural changes on compositional generalization (Csordás et al.,, 2022), explored techniques involving data augmentation (Andreas,, 2020; Cazzaro et al.,, 2024; Chai et al.,, 2023; Jiang et al.,, 2022), injecting structural knowledge via additional supervision or auxiliary tasks (Chakravarthy et al.,, 2022; Jiang and Bansal,, 2021), and investigated the limits of standard transformers on compositional tasks (Dziri et al.,, 2023; Zhou et al., 2023b, ).
After the review, we discuss how these new approaches may inform our understanding of human compositionality, including their implications for its neural mechanisms and development. We conclude with a discussion of the ‘mere implementation’ objection, the second horn of Fodor and Pylyshyn’s (1988) argument.
4.2 Architectural Inductive Biases
A major focus of the work on compositionality in neural networks has been the search for architectural inductive biases that would allow neural networks to generalize on compositional tests like the SCAN task described above. The inductive biases that have been explored vary in the extent to which they explicitly incorporate the exact kind of compositionality characteristic of classical or symbolic systems. Neuro-symbolic hybrid architectures incorporate explicit symbolic operations into neural network models, while strictly neural architectures attempt to implement inductive biases that could plausibly facilitate the emergence of structure-sensitive, compositional processing without the use of such explicit symbolic operations. [^19] The term “neuro-symbolic” is sometimes used to describe both of these kinds of approaches together—those that attempt to integrate explicit symbolic representations or operations in a neural network architecture as well as the approaches that take inspiration from symbolic or classical computation to design strictly neural architectures. However, we believe the distinction between these approaches is important, and therefore reserve the term “neuro-symbolic hybrid” for the former. This distinction has also been usefully described as a difference between “neuro-compositional” and “neuro-symbolic” or between “neurons-in-structure” and “structure-in-neurons” (Smolensky et al., 2022b, ). Each of these approaches is a very active area of research, so a comprehensive review is beyond the scope of this chapter, but in the following we provide a targeted survey (see Yu et al.,, 2023; Smolensky et al., 2022b, ; Webb et al.,, 2024, for other reviews).
Neuro-symbolic hybrids (Ellis et al.,, 2021; Gaur et al.,, 2022; Marcus,, 2020; Marcus and Davis,, 2020; Yu et al.,, 2023; Zhou et al.,, 2024) are systems that attempt to fuse neural and symbolic modules or components into a single system that garners the benefits of each—the power and flexibility demonstrated by modern deep learning systems in dealing with high-dimensional inputs like text and images (LeCun et al.,, 2015), combined with the strength of symbolic modules in ensuring that representations or processes are compositional (Fodor and Pylyshyn,, 1988; Marcus,, 1998).
For example, Neural Module Networks (Andreas et al.,, 2017, 2016; Bahdanau et al.,, 2020) assign separate neural “modules” to specific primitive operations (e.g., filtering or querying an image based on certain attributes like color or shape, or checking if two colors are the same). These modules can be composed like symbolic functions to dynamically construct a network to accomplish a specific series of operations. For example, in the visual-question answering setting, they can be used to answer questions like: “Do the green object to the left of the red sphere and the blue object to the right of the orange cylinder have the same shape?" (Johnson et al.,, 2016). When they are tested in a compositional generalization setting, they show some advantages relative to standard networks, but still appear to be inferior to humans (Bahdanau et al.,, 2018, 2020).
In another neuro-symbolic approach, Neural Program Synthesis (Nye et al., 2020a, ; Xu et al.,, 2019; Zhong et al.,, 2023), building on work in program synthesis (Gulwani et al.,, 2017) and probabilistic program induction (Lake et al.,, 2015), explicit symbolic rules or programs are inferred from data using neural networks. These explicit rules can then be manipulated or processed using traditional programming languages to harness the compositional power of these classical systems. On the SCAN task, for example, a neuro-symbolic model based on neural program synthesis (Nye et al., 2020a, ) was able to infer the correct grammar from only a few examples and generalize to given test samples, thereby performing near-perfectly on the compositional split. In general, neuro-symbolic models have demonstrated an impressive ability to learn visual and linguistic concepts from data, and to generalize these to unseen examples (Ellis et al.,, 2021; Mao et al.,, 2019).
It is important to note that while the neuro-symbolic approach may show that it is possible for neural networks to implement compositional generalization behaviors when they are augmented with explicit symbolic processes, it cannot show that the neural components themselves offer an alternative explanation of these generalization behaviors. In other words, these systems are designed to ‘merely implement’ the same kind of compositionality present in classical symbolic systems, and so do not so much grapple with the second horn of the compositionality challenge (2.3), as embrace it (for discussion, see 5.3).
Work investigating strictly neural architectural inductive biases, on the other hand, seeks to meet this challenge by showing that compositional generalization does not require explicit rules, symbols or hand-coded representations (Altabaa et al.,, 2023; Kriete et al.,, 2013; Palangi et al.,, 2017; Rougier et al.,, 2005; Russin et al., 2020a, ; Soulos et al.,, 2023; Schlag et al.,, 2019; Webb et al.,, 2024, 2023). [^20] Again, this approach has been referred to as “neurocompositional” or “structure-in-neurons,” as opposed to the “neuro-symbolic” or “neurons-in-structure” approach (Smolensky et al., 2022b, ). The goal of this work is to find an architectural inductive bias generic enough that it can be applied across many domains (e.g., language tasks, vision tasks, etc.), while still empowering DNNs to generalize in new ways (Battaglia et al.,, 2018). If such an architectural feature is discovered, it may shed light on how the architecture of the human brain could have evolved a bias for compositionality without building in overt symbolic operations, which may have been unlikely to emerge in a biological system (O’Reilly et al.,, 2014).
Some of this work has taken direct inspiration from symbolic systems, however, and attempted to translate compositional features of classical computing, like role-filler independence (Chakravarthy et al.,, 2022; Palangi et al.,, 2017; Schlag et al.,, 2019; Smolensky,, 1990) and variable binding (Webb et al.,, 2021), into architectural features that could plausibly be encoded in a strictly neural system. Other work has tried to replicate some of the architectural motifs known to exist in the brain, like those associated with the prefrontal cortex (Russin et al., 2020b, ), the parietal cortex (O’Reilly et al.,, 2021; Summerfield et al.,, 2020), or the medial temporal lobe (Behrens et al.,, 2018; Whittington et al.,, 2020).
Figure 4: Architectural inductive biases. Modifying a standard encoder-decoder architecture (Bahdanau et al.,, 2014) can instantiate a different set of inductive biases, which can have important implications for how the model generalizes. One such bias—the separation of representations for structure and content (O’Reilly et al.,, 2021; Russin et al., 2020a, )—can facilitate compositional generalization on the SCAN task, transforming an o.o.d. generalization problem into two separate i.i.d. generalization problems (one involving a known structure, such as the relationship between adverbs and verbs, and one involving known content, such as the meaning of the primitive verb “jump”). Figure adapted from Russin et al., 2020a .
One major theme of this work has been the use of architectural motifs that encourage the model to learn relatively independent representations for structural (e.g., “role”) information and content (e.g., “filler”) information. The idea here is that if structural information (e.g., general knowledge about verbs or adverbs in a given language) is learned separately from content information (e.g., knowledge about a specific verb like “walk”), then this will allow abstract, content-independent structural knowledge to be generalized to other specific contents (e.g., “jump”).
For example, tensor product representations (Smolensky,, 1990), in which vectors representing “role” information and vectors representing “filler” information are bound by the tensor product operation to construct composite representations, has been extended to the deep learning framework in a number of different ways (Jiang et al.,, 2021; McCoy et al.,, 2019; Palangi et al.,, 2017; Schlag et al.,, 2019; Soulos et al.,, 2020, 2022; Smolensky et al., 2022a, ). The original work with tensor products (Smolensky,, 1990) sought to establish how a strictly neural system could embody the principle of compositionality and represent constituent structure (Fodor and Pylyshyn,, 1988), but was unconcerned with the problem of learning such representations from data. However, this newer work explores tensor product representations in the context of learning by building tensor-product-like operations into DNN architectures such as RNNs (Palangi et al.,, 2017) or transformers (Schlag et al.,, 2019), and measuring the extent to which it encourages the models to represent constituent structure in an interpretable, compositional way (Russin et al., 2021a, ; Soulos et al.,, 2020; Smolensky et al., 2022a, ). These networks have proven capable of performing well on tasks such as mathematical word problems (Schlag et al.,, 2019), machine translation (Soulos et al.,, 2022), and abstractive summarization (Jiang et al.,, 2021), though they have been shown to fail in some cases of o.o.d. generalization and extrapolation (e.g., in the math setting; Schlag et al.,, 2019).
Other work has implemented similar ideas about role-filler independence and variable-binding, but without the use of explicit tensor product representations (Chakravarthy et al.,, 2022; Russin et al., 2020a, ; Jiang et al.,, 2024; Webb et al.,, 2024). For example, Russin et al., 2020a showed that when a strict separation between structure-related information and content-related information is imposed in the attention mechanism of an encoder-decoder RNN architecture (Bahdanau et al.,, 2014), performance on the compositional generalization test in the SCAN task is improved to 78.4% accuracy. Concurrent work based on a similar principle showed that this performance can be improved even further to 98.8% when a few small changes, such as adding noise, are introduced (Li et al.,, 2019). The structure-content separation in these models is imposed not by hand-coding specific representations, but by modifying the architecture so that separate information-processing streams are responsible for (1) choosing what to attend in the input at each time step (we can label this the “structure” stream), and (2) for translating the individual instruction-words that are currently attended into their corresponding actions (we can label this the “content” stream). This modification forces the model to “factorize” (Behrens et al.,, 2018; Whittington et al.,, 2020) the problem into two sub-problems: the first is to find the right attentional procedure or routine (i.e., the right sequence of decisions about where to attend at each time point), and the second is to memorize how individual primitive verbs should be translated into their corresponding actions. In effect, this factorization turns an o.o.d. problem into two i.i.d problems (see Figure 4). Rather than confronting a novel structure-content combination that lies outside of its training distribution, the model can combine two in-distribution tasks: one where the individual verb is translated into its corresponding action (i.i.d. with respect to training because “jump” → JUMP appears in the training set), and one where the routine for a given grammatical structure is executed (i.i.d. with respect to the model’s training because all constructions with other primitive verbs appear in the training set).
An alternative way of thinking about this kind of separation is as a “relational bottleneck” (Altabaa et al.,, 2023; Campbell and Cohen,, 2024; Campbell et al.,, 2023; Webb et al.,, 2024): these architectural motifs force the network (or part of the network, e.g., the “role” or “structure” stream) to represent information that is purely relational, or independent of any content associated with individual words or objects. This encourages compositionality because it allows generalization to novel content-relation or content-structure combinations, as long as the individual contents and relations are familiar (i.e., they are i.i.d. with respect to the training data). These ideas have been explored in a number of different settings, including in the Emergent Symbol Binding Network capable of solving a task based on Raven’s Progressive Matrices (Webb et al.,, 2021) and in applications to abstract problem-solving tasks (Altabaa et al.,, 2023).
Both neuro-symbolic hybrids and strictly neural architectural inductive biases have proven successful in exhibiting compositional generalization behaviors in limited domains such as the SCAN dataset and related tasks (Nye et al., 2020b, ; Russin et al., 2020a, ; Ossowski et al.,, 2024). Both approaches can be understood as succeeding on o.o.d. generalization problems by instantiating powerful inductive biases in learning systems: neuro-symbolic models leverage overtly classical elements such as symbolic or rule-like representations, whereas neural architectural inductive biases attempt to realize similar properties in strictly neural systems. These inductive biases can in some cases turn what looks like an o.o.d. generalization problem into multiple separate i.i.d. generalization problems (see Figure 4; Russin et al., 2020a, ).
While both of these approaches are active areas of research, there has been some difficulty in scaling them up to difficult tasks such as language modeling, where large, unbiased transformer architectures trained on huge datasets have dominated (Bubeck et al.,, 2023; Brown et al.,, 2020). This may be because it is more difficult to design strong inductive biases (whether neuro-symbolic or strictly neural) by hand than it is to learn them from data, a familiar dynamic encapsulated in the “bitter lesson” (Sutton,, 2019). The remaining sections of this review discuss work exploring whether such data-driven methods, including metalearning and language-model pretraining, can also endow neural networks with the inductive biases required for compositional generalization.
4.3 In-Context Compositionality via Metalearning
In the metalearning paradigm (Binz et al.,, 2023; Griffiths et al.,, 2019; Sandbrink and Summerfield,, 2024; Wang et al.,, 2018; Wang,, 2020), rather than building inductive biases into a network’s architecture, they are imparted to the model by specifying a distribution of tasks on which it is trained. This has the advantage that the specific architectural features necessary to achieve the learning behaviors of interest do not need to be hand-designed in advance (McCoy et al.,, 2020). Instead, the researcher (who has learned their “bitter lesson”; Sutton,, 2019) can simply define the set of learning problems at which the model should excel and rely on the optimization process during metalearning to discover whatever inductive biases are useful for accomplishing these behaviors. Thus, metalearning neural networks learn how to learn—they are trained to rapidly acquire new skills or generalize in specific ways on new problems.
Figure 5: The inner-loop learning algorithm that emerges in a network’s activation dynamics may have different properties than the original learning algorithm used to train the network’s weights. These two learning algorithms can simultaneously coexist in a single network.
A number of metalearning methods have been developed for use with neural networks (Bengio et al.,, 1991, 1995; Duan et al.,, 2016; Finn et al.,, 2017; Hospedales et al.,, 2020; Schmidhuber,, 1987). In many of these methods, the model learns to implement in its forward activation dynamics a separate “inner-loop” or “in-context” learning (ICL) algorithm, distinct from the “outer-loop” or “in-weight” learning (IWL) algorithm used to train the network in the first place (see Figure 5). In this case, the network’s inner-loop algorithm “learns” within the flow of information from its inputs to its outputs (Duan et al.,, 2016; Lake,, 2019; Lake and Baroni,, 2023; Santoro et al.,, 2016; Wang et al.,, 2018). The network’s weights are trained such that the forward pass of the model can take some study examples as inputs in context and learn from them to produce the correct generalization behavior on a test query (see Figure 6). For example, a sequence-processing network such as an RNN or a transformer can be trained on inputs that consist of (1) a set of study examples along with their labels, and (2) a single test query whose label is not given, but must be inferred. In this case, the network is trained to produce a prediction about the label corresponding to the test query, conditioned on the set of training samples and their labels. The model is then evaluated by freezing its weights and testing on new study examples and a new test query drawn from another task.
Figure 6: Metalearning setting. (A) In a typical (non-metalearning) training scenario, a learner presented with a compositional generalization problem is trained on the training set (green, numbered 1, 2, 3) and then tested on the test set (blue). In this case, compositional generalization requires o.o.d. generalization. (B) In a metalearning scenario, the metalearner is trained on a distribution of tasks (“episodes”). In each task, the metalearner is presented with a set of in-context study examples (green) that are analogous to the training set in the vanilla learner. The metalearner is required to generalize out-of-distribution from these study examples to a test query (blue), which is analogous to the test set in the vanilla learner. By (meta)training on many such o.o.d. generalization problems (red, numbered 1, 2, 3), the metalearner can come to implement in its forward dynamics a separate learning algorithm with an inductive bias for compositionality. When a new o.o.d. generalization problem is sampled from the same distribution of tasks for (meta)testing (purple), the metalearner can generalize because it is familiar with such tasks (i.e., the task is i.i.d. with respect to its training data), even though the problem itself involves o.o.d. generalization.
In this setting, we can distinguish the outer-loop (“base”) learning algorithm that determines the updates to the network’s weights during metalearning, and the inner-loop (“mesa-optimization”; Hubinger et al.,, 2021) or in-context learning algorithm that is implicitly implemented in the network’s forward dynamics (see Figure 5). The inner-loop algorithm does not require updates to the weights of the network to learn (i.e., it can still learn a new task when the weights are frozen).
Crucially, the inner-loop learning algorithm implemented in the model’s forward pass can have different properties than the outer-loop learning algorithm used to train the weights of the model (see Figure 5). For example, the inner-loop learning algorithm may have a completely different learning rate or updating procedure, or embody inductive biases that are not present in the outer-loop algorithm.
This suggests a novel perspective on how a neural network might instantiate compositionality without incorporating explicit classical symbols: by (meta)learning from data to implement in its activation dynamics a learning algorithm with an inductive bias for compositionality. Even if the outer-loop learning algorithm used by neural networks lacks compositionality, as theorists such as Fodor and Pylyshyn, (1988) and Marcus, (1998) suggest, it is still possible that the inner-loop learning algorithm that a network comes to implement in its activation dynamics will exhibit the property of compositionality.
Lake and Baroni, (2023), following up on earlier work (Lake,, 2019), applied this approach to compositional generalization problems such as the SCAN task (Lake and Baroni,, 2018), and showed that a metalearning transformer was capable of near-perfect accuracy on the task. The transformer was trained on a distribution of tasks, each of which was itself a compositional generalization problem. In particular, each task had a few samples from the grammar that were in principle sufficient to infer the rules that could be used to produce the correct answer on the test sample. The model was trained on some generalization tasks, and tested on ones that it had not seen. The results showed that the transformer was capable of metalearning to implement a compositional learning algorithm in its activation dynamics, achieving near-perfect accuracy on held-out compositional generalization tasks that were not seen during metalearning.
These results demonstrate a novel strategy for endowing neural network models with the property of compositionality, and for overcoming the problem of o.o.d. generalization. After metalearning was complete, the model was given an o.o.d. problem in context, and its inner-loop learning algorithm exhibited an ability to generalize. However, from the perspective of the outer-loop learning algorithm, the problem was familiar (i.i.d.), because it was trained on similar o.o.d. generalization tasks during metalearning (see Figure 6). Thus, metalearning creates a scenario where two kinds of generalization can coexist simultaneously: the outer-loop learning algorithm is simply generalizing i.i.d., while the inner-loop learning algorithm is capable of generalizing o.o.d. because of the inductive biases imparted to it by outer-loop learning (see Figure 7).
Figure 7: Outer-loop generalization can be i.i.d., even if each generalization problem solved in the inner-loop is o.o.d. A metalearner can be trained on a distribution of tasks, where each task is itself a generalization problem like the SCAN task (see Figure 3). In this case, each sample (“episode”) in the outer loop corresponds to an o.o.d. generalization problem that must be solved by the inner-loop learning algorithm. The metalearner can be tested on new tasks from the same distribution (i.e., the outer-loop generalization problem is i.i.d.), where each new task is itself an o.o.d. generalization problem. In this figure, the colors are coordinated with Figure 6, where red and purple represent train and test samples in the outer loop, while green and blue represent the train and test samples in the inner loop.
Lake and Baroni, (2023) conclude that although their metalearner could generalize on o.o.d. problems that were similar to those seen during metalearning, it could not generalize to completely new kinds of problems (i.e., ones that were o.o.d. with respect to the metalearning dataset). One way of interpreting this finding is to maintain that the kind of compositionality that humans exhibit is truly o.o.d., so the model’s failure shows that it is inadequate as a model of human compositionality. However, another interpretation invites a reexamination of human compositionality and suggests that it too could be a property of an inner-loop learning algorithm, in this case implemented by the human brain (Russin et al., 2024a, ). This would challenge the claim that humans are capable of generalizing compositionally in completely novel domains, instead establishing the weaker claim that humans generalize o.o.d., but only in specific domains determined either by evolutionary history or by prior learning experiences.
4.4 In-Context Compositionality via Pretraining
The original tasks used to evaluate compositionality (e.g. the SCAN dataset; Lake and Baroni,, 2018) were developed at a time when most models were trained from scratch to perform a specific task, such as machine translation (Bahdanau et al.,, 2014; Vaswani et al.,, 2017). However, researchers soon began to utilize self-supervised methods that did not require human labels (e.g., annotations, translations, etc.), thus allowing models to be trained on the huge amounts of unlabeled text available on the internet. The dominant self-supervised task has been language modeling, where a neural network parameterizes a distribution over sequences of natural language text. For example, in causal language modeling this distribution is expressed in terms of the probability of the next token given the entire preceding context. Researchers found that as transformer models were scaled up (in terms of number of parameters and amount of data) language modeling performance continued to improve. This meant that better performance could be achieved if large language models were pretrained on unlabeled text and then fine-tuned on the specific tasks of interest, rather than training from scratch on the specific task (Devlin et al.,, 2019; Peters et al.,, 2018; Radford et al.,, 2019).
Researchers then discovered that at even larger scales (e.g., at the scale of GPT-3, 175 billion parameters), you could in some cases forgo the fine-tuning step and simply ask these large language models to do a new task in context by providing a few examples of the task to be performed (Brown et al.,, 2020). For example, rather than fine-tuning a pretrained model to predict whether a human-labeled restaurant review was positive or negative, you could prompt the model with a few study examples consisting of reviews and their corresponding labels, and test it on a test query—an unlabeled review (also included in the same context at the end of the prompt). This ability to flexibly solve new tasks that were explicitly instructed or demonstrated with a few examples, termed in-context learning (ICL), was further improved when models were fine-tuned on human-generated instruction-following datasets (Wei et al., 2022a, ) and with reinforcement learning by human feedback (RLHF; Ouyang et al.,, 2022). This created yet another paradigm shift in natural language processing: very large models are trained on very large datasets and then tested in context on specific tasks.
From a theoretical perspective, this “pretrain-test” paradigm (Hupkes et al.,, 2023) has important similarities with the metalearning setting described above. In metalearning, models acquire inductive biases by training on a distribution of tasks. The pretraining step in LLMs, where they are trained to predict the next word on a very large dataset, can be seen as training on a distribution of tasks (see Figure 8; Brown et al.,, 2020; Chan et al.,, 2022; Radford et al.,, 2019). Sometimes this prediction will look more like guessing based on frequencies (e.g., memorizing a common idiom), but other times the task can involve more sophisticated behaviors such as understanding recursion in computer programming (see Figure 8). There are doubtless many instances in these large datasets where the best way to predict the next word is to learn something in context, and perhaps even to compose novel concepts in context (see “in-context composition” in Figure 8). Thus, although it is easy to dismiss next-word prediction as a simplistic or purely statistical training objective, upon further reflection one can see how it might put pressure on the model to develop an inner-loop or ICL algorithm in its forward activation dynamics, and perhaps even to endow this inner-loop algorithm with the property of compositionality.
LLMs at this scale perform quite well on many kinds of tasks that seem to require compositionality. As anyone who has interacted with chatGPT or similar products knows, these models are capable of writing whole paragraphs of coherent text about almost any topic, including novel concepts or made-up words that are included in the prompt. They can construct sentences with sophisticated grammatical structure (Linzen and Baroni,, 2021), and can put together well-reasoned arguments (Herbold et al.,, 2023). They even recapitulate some of the syntactic phenomena thought to require innate grammatical constraints such as syntactic island effects (Wilcox et al.,, 2023, although see Lan et al., 2024). They are also surprisingly adept at computer programming, and are capable of writing novel functions or scripts (e.g., involving compositions of known functions) to accomplish a goal given in natural language instructions. These models have achieved state-of-the-art performance on natural language inference tasks such as recognizing logical entailment (Brown et al.,, 2020), and have even demonstrated human-like tendencies in analogical reasoning tasks such as Raven’s Progressive Matrices (Webb et al.,, 2023), which arguably require compositionality as well.
Researchers have also evaluated these models on the tasks designed specifically to test compositional generalization capabilities (Lake and Baroni,, 2018). As in the metalearning setting, the evaluation works slightly differently than in the original SCAN dataset, where models were trained from scratch on tens of thousands of examples without “jump” and tested on instructions containing “jump.” In the pretrain-test setting, the pretrained LLMs are given just a few study examples from the SCAN task in context and asked to generalize to new test samples that are also given in the same input. The compositional generalization test can be replicated by excluding certain kinds of examples from the context (e.g., longer examples, or more complicated constructions involving “jump”). When GPT-3 (code-davinci-002) was evaluated with a standard prompt on a compositional split of the SCAN task in this way, it only achieved 16.7% accuracy (Zhou et al., 2023a, ). [^21] Other studies have tested different versions of GPT with slightly different setups (e.g., slightly different prompts) and found similar results, although GPT-4 achieved 58% when the study examples given in the prompt were sorted by length (Lake and Baroni,, 2023). However, the inclusion of novel prompting methods improved performance to a near-perfect accuracy of 99.7%. These prompting methods, called “chain-of-thought” prompting (Wei et al.,, 2023) and “least-to-most” prompting (Zhou et al., 2023a, ), have been shown to improve performance on many reasoning tasks. In chain-of-thought prompting (Kojima et al.,, 2023; Wei et al.,, 2023), the model is asked to “think step-by-step,” and/or given examples that show intermediate steps of reasoning. In least-to-most prompting (Drozdov et al.,, 2022; Zhou et al., 2023a, ), these in-context examples are simply given in order of least difficult to most difficult. Many other prompting strategies have been developed to further improve performance on similar tasks (Huang and Chang,, 2023; Nye et al.,, 2021; Press et al.,, 2023; Yao et al.,, 2023).
The fact that the prompting strategy makes such a difference in the resulting measures of performance may be taken to indicate that the compositional generalization exhibited with certain prompts is fragile and cannot reflect a genuine competence (Bender and Koller,, 2020; Bender et al.,, 2021; Lanham et al.,, 2023; Turpin et al.,, 2023; Webson and Pavlick,, 2022). However, it is well understood that human performance depends in large part on the instructions they are given (Evans et al.,, 1994), and the specific content or domain on which their abilities are tested (Klauer et al.,, 2000; Wason,, 1968). LLMs exhibit similar content effects (Dasgupta et al.,, 2023), highlighting the need for fair comparisons between humans and machines (Firestone,, 2020; Buckner, 2023a, ).
Aside from these language-specific compositional generalization settings, much work has focused on evaluating whether large vision models or vision-language models exhibit similar compositional abilities (Ma et al., 2023b, ; Conwell and Ullman,, 2022; Zheng et al.,, 2024). The results here have been mixed, suggesting a dissociation between the performance measured on compositional generalization problems developed in language-specific settings and those developed in other domains.
Figure 8: Predicting the next word on a large corpus of text can be seen as a kind of metalearning (Brown et al.,, 2020; Radford et al.,, 2019), in which the LLM trains on tasks requiring diverse capabilities such as memorization of factual knowledge, idioms, inflectional morphology (past tense), programming, and identifying the indirect object of a sentence (Wang et al.,, 2022). In some cases, successful prediction of the next word may require in-context composition of novel elements. Analogous to the metalearning scenario, a distinction can be made between the outer-loop learning algorithm training the weights of the network, and an implicit inner-loop learning algorithm that operates on information given in context (green) to make inferences relevant to predicting the next word(s) (blue).
Taken together, the impressive performance of LLMs shows that at least in some cases, compositionality can arise as an emergent property of the ICL algorithm implemented in the forward activation dynamics of a large model trained to predict the next word on a large corpus of text (and in some cases also instruction-tuned and trained with reinforcement learning by human feedback). Similar to the studies using metalearning models (Lake,, 2019; Lake and Baroni,, 2023), these LLMs demonstrate that even though the outer-loop learning algorithm may not itself be capable of compositional (o.o.d.) generalization, it can endow an inner-loop, ICL algorithm with this capability. This again shows that the o.o.d. generalization problem solved by the inner-loop ICL algorithm may be an i.i.d. generalization problem from the perspective of the outer-loop learning algorithm, which has probably had many opportunities to learn how to learn similar problems in context from its training set (see Figure 8).
Can these findings about LLMs inform us about human compositionality? It is of course difficult to make any inferences from LLMs to humans because there are so many differences between them. For example, transformers are not biologically plausible in many ways, the scale of the training dataset is orders of magnitude larger than a human lifetime’s worth of language data, etc. (Frank,, 2023; Linzen and Baroni,, 2021; Pavlick,, 2023; Warstadt and Bowman,, 2024). However, these results do show that in principle it is possible to metalearn an ICL algorithm capable of compositional generalization by simple self-supervised learning in the outer-loop. Furthermore, as opposed to the experiments with explicit metalearning (Lake and Baroni,, 2023), the impressive compositional behaviors of LLMs show that the distribution of tasks on which a metalearner is trained doesn’t need to be deliberately contrived to encourage compositionality—strong compositional generalization abilities emerge in these models even when they are trained on unstructured text from the internet. This suggests that humans, who also learn from relatively unstructured experiences, could plausibly metalearn throughout development to implement a compositional inner-loop learning algorithm as well (Russin et al., 2024a, ).
5 Discussion
Classical cognitive theorists like Chomsky, (1965) and Fodor, (1975) developed compelling arguments that compositionality is a central property of human cognition, crucial for explaining the creativity, productivity, and systematicity of language and thought (see section 2.1). Neural network modeling was dismissed as an empirically inadequate explanatory paradigm because these models seem to lack the kind of combinatorial constituent structure intrinsic to classical architectures based on atomic symbols and governed by syntactic rules (see section 2.3; Fodor and Pylyshyn,, 1988).
Even as deep neural networks blew past competing frameworks in virtually every area of AI, initial investigations into their compositionality seemed to confirm traditional intuitions that they were not capable of replicating the kinds of compositional generalization that humans exhibit (see section 4.1; Lake and Baroni,, 2018; Kim and Linzen,, 2020). One way of understanding these initial findings is from the statistical learning perspective: compositional generalization requires out-of-distribution (o.o.d.) generalization or extrapolation, but neural networks assume that train and test data will be independent and identically distributed (i.i.d.). Standard neural networks fail on these problems, then, because they lack the strong inductive biases that would allow them to generalize outside of their training distributions. Some have taken these results to demonstrate that neural networks will never achieve compositionality on their own and concluded that they must therefore be augmented with classical symbolic processes (“neuro-symbolic hybrids,” see section 4.2; Marcus and Davis,, 2020; Marcus,, 2020).
More recent advances complicate this picture, suggesting alternative ways in which strictly neural models might explain the same generalization behaviors (Baroni,, 2020). These advances include (1) architectural inductive biases (section 4.2), where a specialized architecture endows the network with an inductive bias for compositionality, (2) metalearning (section 4.3), where a network is trained on many compositional generalization problems so that it learns to implement a compositional inner-loop learning algorithm, and (3) large pretrained language models (section 4.4), which also implement an emergent compositional inner-loop learning algorithm, but come to do so via large-scale pretraining on next-word prediction.
These methods for imparting inductive biases for compositionality into neural networks overcome the challenge of o.o.d. generalization in different ways. Architectural inductive biases that separate structure-related processing from content-related processing (Russin et al., 2020a, ; Behrens et al.,, 2018; Schlag et al.,, 2019; Webb et al.,, 2024) can be seen as transforming an o.o.d. compositional generalization problem into two i.i.d. generalization problems—one involving structure and the other involving content (see Figure 4; Russin et al., 2020a, ). Factorizing structure from content in this way allows generalization to any novel structure-content combination, as long as the particular structures are familiar. The metalearning and large pretraining settings offer a different perspective on the challenge of o.o.d. generalization. In both of these cases, what looks like an o.o.d. problem for an inner-loop or in-context learning algorithm can be seen as i.i.d. for the outer-loop learning algorithm that trained the network in the first place (see Figures 6 and 7); Lake and Baroni,, 2023).
It can be argued that none of these neural networks are actually generalizing outside of their training distributions, and that they have therefore failed to explain human generalization capabilities. However, another perspective on these results is that they show that the generalization behaviors taken as diagnostic of compositionality in humans are possible without strongly out-of-domain generalization.
More generally, these findings provide novel perspectives on human compositionality, suggesting that architectural inductive biases or metalearning may be important for explaining key aspects of these generalization behaviors in humans. In the following, we briefly discuss how these two perspectives can provide novel empirical predictions about the neural mechanisms and development of human compositionality.
5.1 Implications for Neural Mechanisms
Architectural inductive biases The strictly neural approach to architectural inductive biases assumes that there are important innate (non-classical) organizational features of the architecture of the human brain that encourage compositional generalization (Frankland and Greene,, 2020; O’Reilly et al.,, 2014; Russin et al., 2020b, ). Much of the work investigating such architectural inductive biases has explored principles related to the separation of structure and content (Smolensky et al., 2022a, ; Russin et al., 2020a, ; Frankland et al.,, 2021; Webb et al.,, 2024; Whittington et al.,, 2020).
It has been hypothesized that a similar separation constitutes a key organizational principle of the brain (Behrens et al.,, 2018; O’Reilly et al.,, 2021; Summerfield et al.,, 2020). In the visual system, the ventral visual stream (the “what” pathway) encodes object-related information and the dorsal visual stream (the “where” pathway) encodes spatial information (Ungerleider and Mishkin,, 1982). This organization, which has been sculpted over the course of evolutionary history (Mishkin and Ungerleider,, 1982), may encourage generalization to new structure-content (i.e., location-object) combinations in the visual domain (Pylyshyn,, 1989).
More generally, dorsal areas in the parietal cortex seem to encode structure-related information (Summerfield et al.,, 2020), including spatial (Black and Strub,, 1976; Chechlacz and Humphreys,, 2014; Friedman-Hill et al.,, 1995), temporal (Jazayeri and Shadlen,, 2015), and relational (Chafee,, 2013; Chafee et al.,, 2007) information, while ventral areas in the temporal lobe encode content-related information, including information about objects (Reddy and Kanwisher,, 2006) and semantic features (Visser et al.,, 2010). This gross separation between structure and content may extend into memory-related cortical networks (O’Reilly et al.,, 2021; Ranganath and Ritchey,, 2012) as well as the entorhinal cortex and hippocampus (Behrens et al.,, 2018; Manns and Eichenbaum,, 2006; Park et al.,, 2020; Russin et al., 2021b, ; Whittington et al.,, 2020; Zolfaghar et al.,, 2022), where structure-related information represented in medial entorhinal cortex and content-related information represented in lateral entorhinal cortex are bound in hippocampus (Behrens et al.,, 2018; Manns and Eichenbaum,, 2006). This organization of memory-related areas may facilitate generalization to novel structure-content combinations in memory-related domains (Whittington et al.,, 2020).
In addition to this broad separation between structure and content, there are also thought to be specific mechanisms for encouraging abstraction and content-independent processing in the prefrontal cortex (Badre et al.,, 2010; Frank and Badre,, 2012; Kriete et al.,, 2013; O’Reilly and Frank,, 2006; Rougier et al.,, 2005; Russin et al., 2020b, ; Wallis et al.,, 2001). For example, the prefrontal cortex has specialized gating mechanisms known to be important for working memory (O’Reilly and Frank,, 2006), and which may be involved in the uniquely human capacity for rule-learning and indirection (Rougier et al.,, 2005; Kriete et al.,, 2013). Furthermore, the prefrontal cortex is thought to learn sequential, top-down attentional procedures for accomplishing specific goals (Miller and Cohen,, 2001; O’Reilly et al.,, 2020). If this learning is isolated from the particular objects or contents that are being attended (similar to the isolation in Figure 4), this may encourage content-independent representations and compositional generalization to emerge.
Unlike state-of-the-art neural networks like the transformers used in LLMs, the brain is a highly heterogeneous neural network architecture with many specialized features. Architectural motifs like the broad separation between structure-related and content-related processing and specialized circuits in prefrontal cortex seem like promising candidates for helping to explain human compositionality, but much work remains to be done to establish the precise contributions made by each of these proposed mechanisms to human generalization behaviors.
Metalearning The metalearning approach hypothesizes that human compositionality is a property of an inner-loop, in-context learning algorithm that has been metalearned by an outer-loop, in-weight learning algorithm (Binz et al.,, 2023; Dubey et al.,, 2020; Lake and Baroni,, 2023; Russin et al., 2024a, ; Wang et al.,, 2018). [^22] Some have assumed that metalearning in neural networks is a better model of learning on an evolutionary timescale than within an individual’s lifetime (Binz et al.,, 2023). Here, we assume that metalearning can take place in the brain over the course of an individual’s lifetime, and that in-context learning and in-weight learning can simultaneously exist within a single system. This idea is directly investigated by Russin et al., 2024b . This hypothesis predicts that when the neural mechanisms supporting in-context learning, rather than in-weight learning, are driving behavior, this should facilitate compositional generalization.
Russin et al., 2024b found that this interplay between in-context and in-weight learning in neural network models explained human compositional generalization performance on a task recently studied in humans (Dekker et al.,, 2022). Both LLMs and metalearning neural networks achieved much better compositional generalization performance on the task when in-context learning rather than in-weight learning was responsible for behavior. Furthermore, Russin et al., 2024b found that the interplay between in-context and in-weight learning (Chan et al.,, 2022), which has similarities to a tradeoff between working memory and reinforcement learning observed in humans (Collins and Frank,, 2018; Rac-Lubashevsky et al.,, 2023), also explained the curriculum effects observed in the human study (Dekker et al.,, 2022).
There is further support from cognitive and computational neuroscience for the idea that in-context learning and compositional generalization share a set of underlying neural mechanisms in humans. The prefrontal cortex is important for basic capacities known to be involved in human in-context learning, such as working memory (O’Reilly and Frank,, 2006), top-down attention (Cohen et al.,, 1990), and cognitive control (Miller and Cohen,, 2001), and, as previously mentioned, is thought to be important for compositional generalization processes such as rule learning (Calderon et al.,, 2022; Collins and Frank,, 2013; Frank and Badre,, 2012; Kriete et al.,, 2013), reasoning (Crescentini et al.,, 2011; Goel,, 2007), and processing complex syntax (Thompson-Schill,, 2005). Furthermore, it has been hypothesized that metalearning (specifically, meta-reinforcement learning) is an important aspect of the functioning of the prefrontal cortex (Hattori et al.,, 2023; Wang et al.,, 2018).
Thus, while it may be difficult to directly establish that human compositional generalization behaviors are emergent consequences of an inner-loop, in-context learning algorithm, there is indirect empirical support for the idea that compositional generalization and in-context learning share an underlying set of neural mechanisms in humans.
5.2 Implications for Human Development
Architectural inductive biases When taken seriously as models of human learning, the architectural inductive biases reviewed in section 4.2 generate concrete proposals about the emergence of compositionality over the course of human development. Both neuro-symbolic hybrids (Marcus,, 2020) and strictly neural architectural inductive biases (O’Reilly et al.,, 2014; Russin et al., 2020a, ; Webb et al.,, 2024) assume that humans are not perfectly unbiased learners (blank slates), and that innate constraints will ultimately play an important role in explaining the kinds of generalization behaviors they exhibit. The key difference between the two is that the former assumes innate representations that are guaranteed to compose, while the latter builds in innate architectural motifs, and then relies on learning to discover the most useful representations to accomplish a given objective, allowing compositionality to emerge over the course of training (Bates et al.,, 1996). The strength of the neuro-symbolic hybrid approach (and of any approach to learning that assumes symbolic or compositioanl representational elements; Goodman et al.,, 2015; Piantadosi,, 2021; Quilty-Dunn et al.,, 2023; Yang and Piantadosi,, 2022) is that it provides a better guarantee of compositionality, but the strength of the strictly neural approach is that it allows greater flexibility to learn from data (Donatelli and Koller,, 2023).
The neuro-symbolic hybrid approach predicts that the generalizations humans make are universally constrained by innate representations, such as grammatical principles (Chomsky,, 1965; Lightfoot,, 1989; Montague,, 1970; Pinker,, 1994; Pinker and Bloom,, 1990) or the available logical primitives (Piantadosi et al.,, 2016). The strictly neural approach, on the other hand, predicts that these generalizations will be determined to a large extent by prior learning experiences (Prince and Smolensky,, 2004). For example, if the brain embodies an architectural inductive bias to separate structure-related processing from content-related processing, then it should be more difficult to generalize to completely novel structures than it is to generalize to novel compositions of familiar structures/contents. There has been a great deal of work investigating the degree to which innate constraints or core knowledge contribute to cognitive development (Carey,, 2011; Spelke and Kinzler,, 2007; Spelke,, 2022; Feiman et al.,, 2022), but few studies have sought to directly test fine-grained empirical predictions made by the specific architectural inductive biases shown to improve compositional generalization in modern neural networks. Future work could take current models and investigate where their inductive biases allow them to generalize compositionally and where they fail, comparing these results against human data. Ultimately, both hybrid and strictly neural approaches accept that humans are not maximally unbiased learners, and have specific innate inductive biases that facilitate compositional generalization.
Metalearning The metalearning approach, on the other hand, relaxes the assumption that these inductive biases need to be characterizable in architectural terms at all, and explores how they might themselves be learned from data. As has been articulated clearly in other work (Binz et al.,, 2023; Lake and Baroni,, 2023; McCoy et al.,, 2020; Wang,, 2020), the metalearning approach is agnostic about whether this occurs on an evolutionary timescale or within an individual human lifetime. Many have shied away from interpreting metalearning neural networks as models of human development (although see Russin et al., 2024a, ; Wang,, 2020; Wang et al.,, 2018), instead choosing to emphasize its utility as a methodological tool (Binz et al.,, 2023; McCoy et al.,, 2020).
When taken seriously as models of human development, however, metalearning models generate testable empirical predictions. If human compositionality is the result of metalearning, we should expect older children to show signs of improved compositional generalization abilities. Existing evidence supports this prediction, suggesting that children learn how to learn more efficiently over time (Bergelson,, 2020), and that the specific ability to compose rules improves over the course of development (Piantadosi and Aslin,, 2016; Piantadosi et al.,, 2018).
A more specific prediction of the metalearning hypothesis is that inner-loop, in-context learning, as opposed to outer-loop, in-weight learning, should improve with experience. This is also consistent with existing evidence, which suggests that older children are much more adept at learning and reasoning in context using working memory and other executive functions (Munakata et al.,, 2012). Indeed, executive functions, which are intimately related to our ability to learn in context (Cole et al.,, 2011; Duncan et al.,, 2017; Hampshire et al.,, 2011; Miller and Cohen,, 2001; Rougier et al.,, 2005), emerge especially late in development and continue to mature throughout adolescence and early adulthood (Ferguson et al.,, 2021).
Finally, the metalearning hypothesis emphasizes that compositional generalization abilities are facilitated by exposure to similar generalization problems. This predicts that humans will be more likely to exhibit compositional generalization behaviors with concepts or stimuli that they have previously encountered. The extensive literature on category learning has demonstrated that humans generalize better when categories are congruent (or verbalizable) with familiar features (Ashby and Maddox,, 2011; Flesch et al.,, 2018, 2022; Love et al.,, 2004). It has also been shown that in domains such as motor learning, humans reliably fail to generalize outside of their training distributions on certain tasks (Zhou et al.,, 2017).
These results are consistent with the idea that humans are capable of generalizing out-of-distribution in certain domains where they have been trained to do so (e.g., learning certain kinds of new words or categories), but fail to generalize in completely novel scenarios. The question of whether a given task is truly out-of-distribution with respect to a lifetime of human experience is difficult to establish empirically. Better naturalistic dataset collection, including egocentric video and audio recordings from children (Sullivan et al.,, 2021; Vong et al.,, 2024), and efforts to curate more ecologically/psychologically plausible language datasets (Warstadt and Bowman,, 2024; Warstadt et al., 2023a, ; Warstadt et al., 2023b, ), should facilitate this line of inquiry.
It is important to note that humans learn from experiences that are highly unstructured compared to the training datasets used in most studies on metalearning (e.g., Lake,, 2019; Lake and Baroni,, 2023; Russin et al., 2024b, ), which are contrived by researchers to engender very specific inductive biases on very specific tasks. While humans are trained on structured curricula in formal educational settings, there is evidence that competences related to compositionality such as sensitivity to hierarchical structure in syntax (Chomsky,, 1957) and to geometric features in visual tasks (Dehaene et al.,, 2022) do not depend on such experiences. However, LLMs have shown that simply predicting the next word on a large dataset of unstructured text can engender an inductive bias for compositionality (Brown et al.,, 2020; Bubeck et al.,, 2023; Zhou et al., 2023a, ). These models develop in-context learning algorithms (von Oswald et al.,, 2023; Xie et al.,, 2022) that demonstrate good generalization performance on compositional tasks (Webb et al.,, 2023; Wei et al.,, 2023; Zhou et al., 2023a, ), even though their training sets were not specifically designed to achieve this. These results are impressive, and provide a proof of concept that an inductive bias for compositionality can be learned from unstructured language data.
However, current LLMs are trained on orders of magnitude more data than humans experience in an entire lifetime (Frank,, 2023; Linzen and Baroni,, 2021; Warstadt and Bowman,, 2024), making it unclear whether similar capabilities could emerge in models trained on a more realistic scale. The emergence of in-context learning abilities has been shown to depend on the distributional properties of the training data and on network architecture (Chan et al.,, 2022). Furthermore, architectural inductive biases such as those related to the “relational bottleneck” (Webb et al.,, 2024) have been shown to improve sample efficiency on certain tasks (Webb et al.,, 2021). Thus, perhaps human compositionality is best characterized as an emergent property of an inner-loop, in-context learning algorithm, but the brain embodies important architectural inductive biases that encourage this algorithm to emerge.
5.3 Mere Implementations?
The empirical evidence reviewed up to this point suggests that neural network models may be able to get past the first horn of Fodor and Pylyshyn’s critical dilemma (see section 2.3). We have focused on this first horn because we believe that recent advances represent a crucial development on this front, eliminating a major obstacle that has hindered the acceptance of neural networks as viable models of human cognition. But what about the ‘mere implementation’ objection, the second horn of the compositionality challenge? Is the success of contemporary DNNs on compositional generalization problems a reflection of the fact that they have implemented classical architectures? And if so, does this mean that they are uninformative for cognitive theory?
Though this is a nuanced and challenging issue, certain cases are unambiguous. As mentioned in section 4.2, neuro-symbolic hybrid models are ‘mere implementations’ by design—proponents of this approach treat the behavioral signatures of compositionality in humans as diagnostic of underlying symbolic mechanisms and set out to explicitly build such mechanisms into neural networks. Though the neural components of these networks are utilized to deal with high-dimensional inputs like images, their compositionality is entirely due to their built-in symbolic components. [^23] This reflects a corresponding hypothesis about human cognition—that it is a combination of two distinct systems, one symbolic and one statistical—which has played a key role in, e.g., the past tense debate (Kirov and Cotterell,, 2018; Pinker and Prince,, 1988; Pinker and Ullman,, 2002). This captures something of the spirit of Fodor and Pylyshyn’s original objection (in fact, they allude to this possibility themselves), and turns it into an empirical hypothesis with a clear predictive upshot—explicit implementation of symbolic mechanisms will be a necessary prerequisite for neural networks to succeed on compositional problems. Current results do not seem to favor this hypothesis, as the findings on strictly neural architectures (section 4.2), metalearning (section 4.3), and LLMs (section 4.4) suggest that such explicit symbolic mechanisms may not be necessary. However, it is still unclear whether neuro-symbolic hybrids will win out in the end. If so, it would mean that DNNs yield few new insights into compositionality, though they may still offer alternative perspectives on other aspects of cognition. [^24] Chalmers, (1993) makes this point. Though compositionality is an important explanatory target, it is just one of the phenomena that an adequate theory of mind must account for. He draws an extended analogy between Fodor and Pylyshyn’s dilemma and a behaviorist who argues that, because a cognitivist will need to account for classical conditioning along behaviorist lines, their approach is not a genuine alternative to the mind and cannot yield new insights. In both cases, the point is the same: local agreement about specific cases need not undermine the robust differences between two explanatory paradigms.
However, the possibility that neural networks are implementing classical symbolic operations is invoked to stave off challenges from DNNs, even in the absence of any tangible predictions of this kind. Without clear mechanistic commitments, it becomes harder to assess. The results reviewed here (sections 4.2, 4.3, and 4.4) indicate that even strictly neural models with no built-in symbolic components may be adequate for replicating the behavioral signatures of compositionality. What should we make of this outcome?
One possibility that was widely discussed in the original connectionist debate is eliminativism (Ramsey et al.,, 1990). An eliminative connectionist would argue that the success of strictly neural models demonstrates that Chomsky and Montague’s shift to treating formal syntactic and semantic theories as models of human psychological processing (see section 2) was a grave mistake (Piantadosi,, 2023), and that the ideal language philosophers’ skepticism about this approach was well-founded. If DNNs can achieve compositional generalization without internally representing any rule-based, combinatorial systems, this undermines the motivation for positing them to explain human behavior. Chomsky was right that this capacity cries out for explanation, but the eliminativist holds that the neural network vocabulary of nodes, weights, layers, attention heads, and so forth, will furnish sufficient explanatory resources for doing so.
The possibility that these networks are implementing classical architectures serves as an obstacle to such eliminativist conclusions (McGrath et al., 2023a, ; McGrath and Russin,, 2024). Physical systems can exhibit functional organization at different levels of description. The fact that no explicit symbols or syntactic operations were built into a transformer that has metalearned to generalize compositionally doesn’t mean that it hasn’t learned to implicitly implement these classical constructs through training. For all we know—runs this version of the objection—the transformers that succeed on these compositional tasks are not alternative cognitive models at all, but mere implementations of the very same theoretical constructs that classical theorists have emphasized all along. On this view, they may help us to understand how classical architectures are implemented in the human brain (McLaughlin,, 1993), but will not be informative at the cognitive level.
Though this is an important empirical possibility, there is a crucial gap between the claim that a pretrained LLM could, for all we know, be merely implementing a classical symbolic architecture and any claim that it must be. The problem with Fodor and Pylyshyn’s dilemma is that it treats mere implementation as a virtual guarantee—if neural networks ever overcome the first horn and exhibit the behavioral signatures of compositionality, it must be because they have merely implemented classical architectures (thereby falling on the second horn). There may be arguments that narrow this gap, [^25] One way to bridge the gap between these two claims would be to argue that the “possibility space” (Cao and Yamins, 2021b, ) is narrow, so that DNNs are likely to be hitting on the very same solutions postulated by classical theorists. One issue with this approach is how to justify the claim that the space of possible solutions is narrow—reliance on intuition here may reflect a failure of contemporary imagination, rather than any genuine limitation. Moreover, even if the solution set is limited, it need not be limited to a single possibility, leaving room for interesting and informative variation between the solutions devised by researchers and those identified by the models themselves. but Fodor and Pylyshyn’s blithe confidence is unwarranted.
The only way to guarantee that the success of DNNs reflects mere implementation would be to treat the behavioral signatures of compositionality as constitutive, rather than diagnostic, of an underlying classical architecture (McGrath et al., 2023b, ). What it is to be a classical architecture, on this constitutive view, just is to be the kind of system that generates this behavioral profile. Adopting this position would ensure that any successful DNN would count as a mere implementation of a classical architecture, but it would do so at the cost of sapping classical theory of the kind of explanatory value that it was originally intended to contribute. The rise of cognitivism and its attendant rejection of behaviorism hinged on the claim that psychological theory must do more than merely redescribe behavior—we need an understanding of the independently identifiable underlying mechanisms that produce the given behavior (Cao and Yamins, 2021a, ). If giving rise to these generalization behaviors is constitutive of being a classical architecture, however, then we preclude even the logical possibility that a DNN could replicate the behavioral signatures of compositionality without implementing a classical system. This means, in effect, that we have failed to specify any theoretical content beyond this characterization of behavior and have lapsed back into mere redescription.
Rather than simply assuming that contemporary DNNs implement classical architectures or searching for some a priori guarantee, what we need to do is try to find out whether or not they are doing so. The burgeoning field of “mechanistic interpretability” is making this increasingly feasible (Elhage et al.,, 2021; Olsson et al.,, 2022; Merullo et al.,, 2024; Wu et al.,, 2023). Though interpretability work is often raised in the context of AI safety considerations (Rudner and Toner,, 2021), it can also play a crucial role in informing cognitive theory, allowing us to extract higher-level, functionally relevant descriptions of these models’ internal representations that we can put into contact with existing cognitive theory (McGrath et al., 2023b, ; Millière,, 2024; Millière and Buckner, 2024b, ). In principle, this process could point us towards eliminativism, if successful models’ representations bear no resemblance to classical constructs, or mere implementationism, if they exactly recapitulate the structure of a fully specified symbolic theory. In the cases where this work has been pursued, however, it seems likely that these models land somewhere in between, recapitulating certain aspects of classical theory, but diverging in interesting ways (Baroni,, 2022; Chen et al.,, 2024; Manning et al.,, 2020; McGrath et al., 2023a, ). In such cases, we have suggested that it may be worth preserving key terms to encompass these behaviors (McGrath et al., 2023a, ; McGrath and Russin,, 2024), as long as it does not lose track of the fact that they contain new insights. In a slogan, not all systems that we may count as implementations are ‘mere’—informative implementations can point us to new insights into previously postulated structures (Smolensky,, 1988; Smolensky and Legendre,, 2011). Though interpretive work specifically focused on evaluating the compositional models reviewed here is in its initial stages, it is a promising line of future investigation (Andreas,, 2019; Lepori et al.,, 2023; Lewis et al.,, 2023; McCoy et al.,, 2019; Russin et al., 2021a, ; Soulos et al.,, 2020; Todd et al.,, 2024). From a philosophical perspective, further progress on deeper issues about the nature of explanation and implementation will put us in a position to use these findings to resolve the mere implementation objection once and for all.
5.4 Conclusion
In this article, we have traced the notion of compositionality from its emergence in early analytic philosophy to its present status as a key explanatory construct in contemporary cognitive science. We have emphasized the role that compositionality plays in debates over the explanatory adequacy of neural network models of human cognition and reviewed a range of recent results suggesting that current deep neural networks are capable of replicating some of the behavioral signatures that have motivated the ascription of internal compositional systems to human beings. In particular, these recent results suggest that architectural inductive biases and metalearning (or large-scale pretraining) are viable approaches to reproducing these behavioral signatures with strictly neural architectures. We have discussed how, when taken seriously as approaches to modeling human compositionality, these frameworks generate concrete predictions about its underlying neural mechanisms and about its development. This recent success also reopens key philosophical questions about the theoretical significance and proper interpretation of neural networks—whether they should be understood as merely implementing classical compositional mechanisms, or may generate novel insights and motivate revisions to classical theories. These pressing questions take us beyond asking what neural networks can do—instead asking what they mean for our understanding of the human mind. We expect that resolving them will require an interdisciplinary approach and increased philosophical engagement.
5.5 Acknowledgments
We would like to thank Richard Kimberley Heck, Josh Schechter, Christopher Hill, Elizabeth Miller, Roman Feiman, Ellie Pavlick, Michael J. Frank, Randall O’Reilly, Edouard Machery, Suraj Anand, Steven Frankland, Taylor Webb, Stavros Orfeas Zormpalas, and the participants in Brown University Dissertation Workshop and the Philosophy and Cognitive Science of Deep Learning reading group for helpful comments and discussions. JR was supported by NIH NIGMS COBRE grant GR5271961.